### v1

# Nanorod Pair Complexes Manipulated via Magnetic Casimir Forces

Abstract
Controlling nanoscale interactions to suppress aggregation from short-range attractive forces is a key problem in nanoengineering. Here, we demonstrate a route to modulate Casmir-Lifshitz interactions between anisotropic nanoparticles with the magnetic fluids. By semi-classical quantum electrodynamics, we study ground state dispersion forces for cylindrical dielectric nanorods made of polystyrene (PS), and zinc oxide (ZnO) embedded in toluene-based host media with gold-coated magnetite nanoparticles and also predict magnetic contributions to the non-retarded excited state interaction. The variation in magnetic permeability enables tuning between repulsive and attractive interaction and a thermally unstable and measurable magnetic Casimir traps are predicted between a pair of ZnO–PS nanoparticles whose equilibrium position can be modulated over an order of magnitude with a small variation in the size of the magnetite nanoparticle. This provides an alternative magnetic Casimir-effect pathway to reversibly tune quantum electromagnetic forces at the nanoscale for assembly and enhancement of colloidal stability.
What if the very surrounding media in space itself could be engineered into a tool for nano-manipulation? For instance, can we take the background media as a magnetic fluid, in which electromagnetic fluctuations exert forces that can pull, push or most interestingly trap matters at the nanoscale level? Such fluctuation-induced interactions, which are referred to as van der Waals 1, 2 (vdW) and Casimir-Lifshitz (CL) 3, 4 forces together, typically act to bring microscopic bodies into contact, leading to unwanted self-aggregation in nanofabrication and soft-matter systems 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15. An important practical issue in nano and micro-devices is to overcome close contact attractions that causes building blocks to cluster together and an ideal solution would be to engineer systems where the interaction turns repulsive at small distances14. Beyond simply mitigating stiction, these fluctuations could also be harnessed to produce stable and controllable interaction potentials. Fundamental investigations of the influence on CL forces by a magnetic surrounding media have been largely restricted to simple geometries 16, 17, 18. Exploiting such a foundation, we have recently expanded the theoretical approach by including the contributions of transverse magnetic (TM) and transverse electric (TE) modes within a model described by the proximity force theorem 19, 20 in such a way that both energy contributions due to TM and TE-modes would partially act as candidates to enable quantum trapping between two quasi-planar surfaces 21.
In this letter, we go beyond these simple geometries and study an experimentally feasible setup which comprises of pairs of cylindrical nanorods made of polystyrene (PS) and zinc oxide (ZnO). The use of the full cylindrical geometry permits a more realistic description of nanoscale behavior, albeit at the price of considerable added theoretical complexity. This stimulates further development of analytical and numerical studies of Lifshitz and vdW interaction between elongated nanoparticles 22, 23, 24, 25, 26, 27. One of us previously derived a few asymptotic limits for this kind of a system 25. Here, we examine the magnetic contributions associated with a surrounding ferrofluid for the long-range and high-temperature limits of ground-state dispersion interactions between cylindrical nanoparticles. We also predict a new kind of magnetic contribution to the retarded excited-state interaction. In an illustrative example, we study the CL interaction between a ZnO nanorod and a PS nanorod suspended in a toluene and gold-coated magnetite nanoparticle (5-7%) fluid mixtures. In the case of two unequal nanorods, we find that subtle change of nanoparticle concentration into the ferrofluid can induce a transition from short-range repulsion to long-range attraction between them. It is the fluid’s magnetic permeability and its dependence upon the magnetite nanoparticle size that dictates both interaction strength and sign, in contrast to the fluid dielectric function which remains nearly constant. Such a device could enable controlled, magnetic-induced entropic CL trapping. Early studies of CL interactions between magnetic fluida 17, 18 were mainly concentrated on the observation of a repulsive force by changing the optical properties of the medium. By contrast, we have implemented here a way in which this zero-frequency, purely entropically driven Casimir trapping is realized exclusively by tuning the fluid’s magnetic permittivity, an effect not investigated earlier.
To lay foundations for such a framework, we develop, within the semi-classical quantum electrodynamics, the general theory for ground and excited state dispersion interactions for dielectric cylindrical nanorods embedded in a magnetic fluid. In the subsequent sections, we present analytical results for the excited state interaction between thin cylinders and the corresponding free energy, followed by numerical analysis of the ground-state forces. We briefly describe the essential optical and magnetic properties of the interacting nanorods and background fluid in the main text and at length in the Supplementary Information. Lastly we demonstrate here that the retardation substantially diminishes the overall interaction strength as well as amplifies the relative weight of magnetic response from the long-range, classical entropic contributions.
Here, we thus propose a novel experimentally feasible approach to manipulate nanorod pair complexes by suspending them in a magnetically diluted fluid, where the interplay between repulsive non-magnetic and attractive magnetic contributions to the CL force can be continuously tuned to form stable magnetic CL traps appearing as equilibria whose strength and range derive not from mechanical considerations but from spectral properties of quantum electromagnetic fluctuations.
Figure 1: Color online: Two cylindrical nanorods, one composed of ZnO with permittivity $\varepsilon_{1}$ and permeability $\mu_{1}$, and the other of PS with permittivity $\varepsilon_{2}$ and permeability $\mu_{2}$, each having an identical radius $a$, are immersed in a magnetic liquid. The magnetic liquid is composed of toluene and it contains dispersed magnetite nanoparticles of diameter $D$ and volume fraction $\phi$ and it has effective permittivity $\varepsilon_{3}$ and permeability $\mu_{3}$. The nanorods are separated by a center-to-center distance $R$.
We consider a system of two parallel cylindrical objects of radii $a$ with dielectric constant and magnetic permittivity, ($\varepsilon_{1}$, $\mu_{1}$) and ($\varepsilon_{2}$, $\mu_{2}$), respectively, embedded in a ferrofluid of dielectric constant $\varepsilon_{3}$ and permittivity $\mu_{3}$ shown in Fig. (1). The cylinders are separated by a center-to-center distance $R$. We assume that the cylinders are infinitely long in the $z$ direction, and we will work in the quasi-static limit where the wavelength of light is much larger than any other length scale in the problem. The dispersion interaction in this system is obtained using scattering approach 25, where the interaction energy per unit length (for a cylinder with length $L$) in finite temperature $T$ is (an explicit derivation is given in the Supplementary Information),
$G(a,R,T)=\frac{k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\,\ln\mathscr{D}(i\xi_{p})$ (1)
where the prime in the summation means that the term for which $p=0$ must be weighted by $1/2$. The above expression is given in Matsubara frequencies $\xi_{p}=2\pi k_{B}Tp/\hbar$, where $p$ is an integer. Here, $\mathscr{D}(i\xi_{p})$ is expressed as the determinant of the scattering matrices found in the “thin cylinder approximation” e.g., the separation distance between the cylinders is much larger than their radii ($R\gg a$). We find that (see Supplementary Information for details),
$\ln\mathscr{D}(i\xi_{p})=f_{0}K^{2}_{0}(\nu_{3}R)+f_{1}K^{2}_{1}(\nu_{3}R)+f_{2}\Big(K^{2}_{0}(\nu_{3}R)+K^{2}_{2}(\nu_{3}R)\Big)$ (2)
where $K_{0},K_{1}$ and $K_{2}$ are modified Bessel functions of zeroth, first and second orders respectively, $\nu^{2}_{\alpha}(i\xi_{p})=\Big(\frac{\xi^{2}_{p}}{c^{2}}\mu_{\alpha}(i\xi_{p})\varepsilon_{\alpha}(i\xi_{p})+k^{2}\Big)$ for $\alpha=1,2,3$, and the coefficients are as follows:
$f_{0}=-\frac{a^{4}}{4}\Bigg[\frac{(\varepsilon_{3}-\varepsilon_{1})(\varepsilon_{3}-\varepsilon_{2})}{\varepsilon^{2}_{3}}+\frac{(\mu_{3}-\mu_{1})(\mu_{3}-\mu_{2})}{\mu^{2}_{3}}\Bigg]\nu^{4}_{3}$ (3)
$f_{1}=-a^{4}\Bigg\{\frac{\nu^{2}_{1}\nu^{4}_{3}}{(\varepsilon_{3}+\varepsilon_{1})(\mu_{3}+\mu_{1})}\Bigg[\frac{(\varepsilon_{3}-\varepsilon_{2})}{\varepsilon_{3}}\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{1}(\mu_{1}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{1})}{2\nu^{2}_{1}}\Bigg)+\frac{(\mu_{3}-\mu_{2})}{\mu_{3}}$ (4)
$\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{1}+\mu_{3})+\varepsilon_{1}(\mu_{1}-\mu_{3})}{2\nu^{2}_{1}}\Bigg)\Bigg]+\frac{\nu^{2}_{2}\nu^{4}_{3}}{(\varepsilon_{3}+\varepsilon_{2})(\mu_{3}+\mu_{2})}\Bigg[\frac{(\varepsilon_{3}-\varepsilon_{1})}{\varepsilon_{3}}$
$\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{2}(\mu_{2}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{2})}{2\nu^{2}_{2}}\Bigg)+\frac{(\mu_{3}-\mu_{1})}{\mu_{3}}\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{2}+\mu_{3})+\varepsilon_{2}(\mu_{2}-\mu_{3})}{2\nu^{2}_{2}}\Bigg)\Bigg]\Bigg\}$
$f_{2}=-\frac{a^{4}\nu^{2}_{1}\nu^{2}_{2}\nu^{4}_{3}}{(\varepsilon_{3}+\varepsilon_{1})(\varepsilon_{3}+\varepsilon_{2})(\mu_{3}+\mu_{1})(\mu_{3}+\mu_{2})}\Bigg[\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{1}(\mu_{1}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{1})}{2\nu^{2}_{1}}\Bigg)$ (5)
$\times\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{2}(\mu_{2}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{2})}{2\nu^{2}_{2}}\Bigg)+\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{1}}\Bigg)\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{2}\mu_{2}}{\nu^{2}_{2}}\Bigg)$
$+\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{2}}\Bigg)\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{1}\mu_{1}}{\nu^{2}_{1}}\Bigg)+\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{1}+\mu_{3})+\varepsilon_{1}(\mu_{1}-\mu_{3})}{2\nu^{2}_{1}}\Bigg)\times$
$\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{2}+\mu_{3})+\varepsilon_{2}(\mu_{2}-\mu_{3})}{2\nu^{2}_{2}}\Bigg)\Bigg]$
where $c$ is the speed of light in vacuum.
Here we are also interested if a dispersion interaction in excited state could be realized between two equal cylindrical nanoparticles. In order to gain an understanding into this issue, it is instructive to start by considering the corresponding analogous situation of two identical atoms; one in its excited state while the other in its ground state. In such a scenario, the interaction can be described as a superposition of a symmetric and an antisymmetric state 28. Among these, the only state that is of practical interest will be the antisymmetric one which is often very long-lived thus effectively holding the system as a trapped metastable state. As a result, a strong resonant dispersion interaction will be maintained between the atoms until (I) the atoms move sufficiently far apart or (II) a photon is emitted, which reduces most of the coherence between them. The interaction energy is due to the separation dependence of the excited energy, such that $E^{\text{res}}(R)=\hbar[\omega_{r}(R)-\omega_{r}(\infty)]$, where $\omega_{r}(R)$ represents the resonance frequency at interatomic separation $R$. The interaction energy of the excited state is 29, 30, $E^{\text{res}}(R)=s^{2}T(R|\omega_{j})\propto R^{-3},$ where $s$ is the magnitude of the transition dipole moment of the excitation and $T(R|\omega_{j})$ is the dipole–dipole coupling tensor (from the electromagnetic Green’s function) at frequency $\omega_{j}$ which accounts for how field from one oscillating dipole influences another at distance $R$ and it typically scales as $R^{-3}$ in the non-retarded limit. This can be compared with the ground state van der Waals interaction which decays with distance as $R^{-6}$31.
This analogy implies that two identical cylindrical nanoparticles can similarly form a long lasting antisymmetric excited state by virtue of a long-range resonant dispersion interaction. A corresponding expression for the retarded excited-state interaction energy, obtained semiclassically in the Supplementary Information, reads
$G^{\text{res}}(a,R,T)$ $=-\frac{2k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\,(t_{0}+t_{1}+t_{2}+t_{3})\,K_{0}(\nu_{3}R)$ (6)
where
$t_{0}=\Bigg[\frac{\varepsilon_{3}-\mu_{1}}{2\,\varepsilon_{3}}\Bigg]a^{2}\nu^{2}_{3};\quad\quad t_{1}=\Bigg[\frac{\mu_{3}-\varepsilon_{1}}{2\,\mu_{3}}\Bigg]a^{2}\nu^{2}_{3}$ (7)
$t_{3}=\Bigg[-\frac{\varepsilon_{1}(\mu_{1}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{1})}{\nu^{2}_{1}}+\frac{2\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}\Bigg]\frac{a^{2}\nu^{2}_{1}\nu^{2}_{3}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})}$
$t_{4}=\Bigg[-\frac{\varepsilon_{3}(\mu_{1}+\mu_{3})+\varepsilon_{1}(\mu_{1}-\mu_{3})}{\nu^{2}_{1}}+\frac{2\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}\Bigg]\frac{a^{2}\nu^{2}_{1}\nu^{2}_{3}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})}$
The entropic contribution ($p=0$) suggests that Eq. (S. 63) provides $R^{-3}$ dependence that is much longer ranged than the ground state interaction between two equal thin cylindrical nanoparticles that is proportional to $R^{-5}$32. It gives both magnetic and non-magnetic contributions. Notably, this previously neglected magnetic contribution in Eq. (S. 63) has the same power law as its non-magnetic contribution. However, the significance of this contribution should be judged by the low temperature/short distance non-magnetic excited state interaction. The non-retarded excited state interaction becomes attenuated by retardation due to the finite velocity of light. In the retarded limit additional effects potentially impact the interaction via real photon exchange. Ultimately, in the limit of large distances or high temperatures, the interaction is expected to be dominated by the non-magnetic and magnetic zero frequency contributions. However, it is important to emphasize that when the interaction occurs in a fluid the excited state lifetime will be significantly reduced. This will impact the potential existence of excited interactions in a fluid at large separations.
The nanoparticles we consider have large enough dimensions to approximate their effective dielectric functions with those of bulk materials. The dielectric functions at imaginary Matsubara frequencies are obtained through the Kronig–Kramer relation
$\varepsilon(i\xi_{p})=1+\frac{2}{\pi}\int_{0}^{\infty}d\omega\,\frac{\omega\,\varepsilon^{\prime\prime}(\omega)}{\omega^{2}+\xi_{p}^{2}}.$ (8)
We use available experimental data from the literature for polystyrene (PS) 33 nanoparticles. This reference is also used to provide the dielectric function for Toluene. For zinc oxide (ZnO) 34 nanorods we use results from density functional theory (DFT) that, together with additional materials information, are described in detail in the Supplementary Information. The optical properties of the magnetic fluid are estimated by taking into account the volume fraction of commercially available gold-coated magnetite particles in a toluene mixture. We use the mixing model presented by Chen et al.35 as,
$\varepsilon_{3}=\frac{\varepsilon_{3,3}(1+2\Lambda_{3})}{1-\Lambda_{3}};\quad\Lambda_{3}=(\phi_{1}+\phi_{2})\frac{\varepsilon_{s}-\varepsilon_{3,3}}{\varepsilon_{s}+2\varepsilon_{3,3}},$ (9)
where the effective dielectric function of the sphere ($\varepsilon_{s}$) is
$\varepsilon_{s}=\frac{\varepsilon_{3,1}(1-2G)}{1+2G};\quad G=\frac{\varepsilon_{3,1}-\varepsilon_{3,2}}{2\varepsilon_{3,1}+\varepsilon_{3,2}}\frac{\phi_{2}}{\phi_{1}+\phi_{2}}.$ (10)
Here, $\phi_{i}$ and $\varepsilon_{3,i}$ are the volume fraction and dielectric function of toluene ($i=3$), magnetite ($i=2$) and gold ($i=1$) with gold coating thickness $R_{\rm Au}$. In the current work, we consider the case where the amount of magnetite is 5% and 7% ($\phi_{2}=0.05,0.07$). Hence, it is of interest to link $\phi_{1}=\phi_{2}\times[(R_{1}/D)^{3}-1]$ where $R_{1}=D+R_{\rm Au}$. We consider four typical cases where the average diameter ($D$) of the magnetite nanoparticles is 10 nm, 13 nm, 15 nm, and 17 nm. We use experimental data from the literature for toluene 33 at $T=298.15$ K. Gold coatings are taken to have bulk properties given in the literature 36. In the main text, we report some important results that are obtained for pure magnetite particles (i.e. for $R_{1}=D$). Most of the results obtained using gold-coating are included in Sec. (S.7) of the Supplementary Information. The size of the magnetite nanoparticles and gold-coated magnetite nanoparticle can be adjusted to control trapping behaviour for pairs of unequal nanorods.
[…]
Figure 2: (Color online): The parametrized dielectric functions on the imaginary axis as a function of the frequencies ${\xi}_{p}$ for PS, and magnetic toluene based on optical experiments given by van Zwol et al.33. For ZnO, we performed DFT calculation [see Sec. (S.5.1) in the Supplementary Information]. The dielectric function for the magnetic fluid is calculated using Eq. (9) and data from Ref. 36 with a volume fraction of $\phi=5\%$.
For water (close to being a non-magnetic fluid) with a fraction $\phi$ of magnetite Klimchitskaya et al. estimated the permeability to be 17, 18
$\mu_{3}(0)=1+\frac{2\pi^{2}\phi}{9}\frac{M_{s}^{2}D^{3}}{k_{B}T}.$ (11)
Toluene is only a little more magnetic ($\mu=0.99996\approx 1.00$) so within the level of accuracy in our work the permeability of both water and toluene-based mixtures can be estimated with the approximation used by Klimchitskaya and Mostepanenko for water-based mixtures 17, 18. Using Eq. (11) we derive the permeability for a range of volume fractions (from 0.05 to 0.07) and diameters (from 10 to 17 nm). The cylindrical nanoparticles we consider are non-magnetic ($\mu_{1}(0)\approx\mu_{2}(0)\approx 1$).
Figure 3: (Color online:) The CL force per unit length between two cylindrical objects (ZnO and PS respectively) with the same radius $a=20$ nm as a function of center-to-center distance $R$ when immersed in a ferrofluid with 5% and 7% concentrations of magnetite nanoparticles of diameter 10, 13, 15 and 17 nm.
We recently investigated the interaction between a planar PTFE surface and a spherical PS surface in the presence of a magnetic fluid 21. In this case both the non-retarded and retarded interaction could be tuned to be repulsive while zero frequency contributions to the CL interaction (both TM and TE) where attractive. This led us to conclude that it is possible to have a magnetic contribution to CL trapping 21. However, trapping still occurred when the magnetic effects were ignored (i.e. the magnetic permeability set to one).
Figure 4: (Color online): The CL force per unit length between two ZnO and PS cylindrical objects was evaluated for volume fraction $\phi=5\%$ and a nanoparticle diameter $D=17$ nm. The terms “without gold” and “with gold” refer to magnetite nanoparticles without and with a gold coating, respectively. For gold coating, layer thickness was taken to be, $R_{\rm Au}=2$ nm. Total1 is the fully retarded calculation for all frequencies without gold coating whereas Total2 reflects the retarded force with gold coating and $p=0$ denotes the zero frequency contribution.
We will in the current work explore a more versatile system consisting of nanorod pair complexes where the materials are described by the dielectric functions shown in Fig. (2). Here the intermediate magnetic fluid has a dielectric function in between those of ZnO and PS. This indicates that the non-retarded and retarded CL interaction between a ZnO and PS nanorod will be repulsive 4.
In Fig. (3) we show that in the case of a pair of ZnO and PS nanorods in a magnetic fluid with toluene mixed with 5-7% magnetite it is possible to obtain trapping that is entirely controlled by the magnetic contribution to the CL force. When the magnetic permeability of the fluid is set to one ($\mu_{3}=1$), the system only experiences repulsion. This dependence of trapping on the magnetic permeability has not been possible to achieve in the past. Notably, when the diameter of the magnetite nanoparticles decreases with fixed volume fraction of magnetite, or a weak magnetic field is applied, leading to reduced magnetic permeability of the fluid 37, the trapping occurs at increasingly larger nanorod pair separations. Ultimately, in the absence of magnetic permeability effects, there is no trapping at all and the nanorod complex is split into two separate nanorods.
Replacing the magnetite nanoparticles with commercially available gold-coated magnetite nanoparticles leads to an additional mechanism to tune the trapping distance as can be seen in Fig. (4). With changes in the gold coating the dielectric function of the background media changes and the trapping distance can be shifted.
This framework suggests a new route for actively tuning the interacting forces between two unequal nanorods where a purely attractive potential can be driven into repulsive–attractive trapping potentials and even into regimes of complete repulsion. By a careful choice of the material parameters, one can tailor finite nanosystems that give rise to reversible adhesive behavior, stable tunable trapping and on-demand destabilization of clusters. These results demonstrate a general platform for nanocluster design in magnetic colloids where adjustable magnetic Casimir forces can be exploited for NEMS/MEMS applications.
Note that the proof-of-principle theory presented here assumes that trapping takes place at separations much larger than the diameter of the nanorod. In this situation, the trapping minima are very weak and can be affected by thermal noise [see Sec. (S.6) in the Supplementary Information]. Nevertheless, this restriction can be mitigated by applying the proximity force approximation (PFA), valid in the short-distance regime, where enhanced stability of the trapping potential is expected.
Going forward, the framework we have developed here indicates some exciting prospects for future extension of tunable dispersion-force control to biomolecular materials. In particular, these ideas could be extended to direct biopolymer complexation with artificial nanothreads via magnetic nanoparticles in biofluids with a suitable choice of materials. This has a great potential in a variety of future diagnostic and therapeutic nanomedicine research 38. It also provides a route to field-controlled assembly in hybrid soft–magnetic nanostructures and could also find utility for biosensing, targeted delivery, reconfigurable nanoarchitectures, and beyond.
1 Acknowledgement
SP and UDG acknowledge support from the Marie Skłodowska-Curie Doctoral Network TIMES, grant No. 101118915 and SPARKLE grant No. 101169225. MB’s contributions to this research are part of the project No. 2022/47/P/ST3/01236 co-funded by the National Science Centre and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 945339. Institutional and infrastructural support for the ENSEMBLE3 Centre of Excellence was provided through the ENSEMBLE3 project (MAB/2020/14) delivered within the Foundation for Polish Science International Research Agenda Programme and co-financed by the European Regional Development Fund and the Horizon 2020 Teaming for Excellence initiative (Grant Agreement No. 857543), as well as the Ministry of Education and Science initiative “Support for Centres of Excellence in Poland under Horizon 2020” (MEiN/2023/DIR/3797). We gratefully acknowledge Poland’s high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016228 and for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through grant no. PLL/2023/4/016319. CP acknowledges the access to high-performance computing resources from NAISS, provided by NSC and PDC. LMW acknowledges financial support from the US Department of Energy under Grant No. DE-FG02-06ER46297.
Supplementary Information
Nanorod Pair Complexes Manipulated via Magnetic Casimir Forces
S Pal , L. M. Woods , C. Persson , I. Brevik ,
U. De Giovannini ,  and M. Boström
*Corresponding authors: palsubhojit429@gmail.com, umberto.degiovannini@unipa.it, mathias.bostrom@ensemble3.eu
S.1 Maxwell’s equations
We start with Maxwell’s equations where there is no current density fluctuations ($J=0$) and charge density fluctuations $(\rho=0)$ written as 39
$\nabla\cdot E=0\,\quad\quad\quad\nabla\cdot D=0\,(D=\varepsilon E)$ (S. 1)
$\nabla\cdot B=0\,\,\,\quad\quad\quad\nabla\cdot B=0\,(B=\mu H)$
$\nabla\times E=-\frac{\partial B}{\partial t}\quad\quad\quad\nabla\times E=-\frac{1}{c}\frac{\partial B}{\partial t}$
$\nabla\times B=\mu_{0}\varepsilon_{0}\frac{\partial E}{\partial t}\quad\quad\quad\nabla\times H=\frac{1}{c}\frac{\partial D}{\partial t}$
We can deduce a further equation from the last two expressions of the right side panel as,
$\nabla\times(\nabla\times E)=-\frac{\mu}{c^{2}}\frac{\partial}{\partial t}(\nabla\times H)\implies\nabla^{2}E+\frac{\mu\varepsilon}{c^{2}}\frac{\partial^{2}E}{\partial t^{2}}=0$ (S. 2)
Throughout the calculations, we consider the time dependency of the parameters $\sim e^{-i\omega t}$. After Fourier transform we obtain an expression of electric field $E$ and magnetic field $H$
$\nabla^{2}E+\frac{\omega^{2}}{c^{2}}\mu\varepsilon E=0$ (S. 3)
$\nabla^{2}H+\frac{\omega^{2}}{c^{2}}\mu\varepsilon H=0$ (S. 4)
We proceed to consider an ansatz of Eq. (S. 3) and of (S. 4) as
$E=E(r,\theta)\,e^{i(kz-\omega t)}\quad H=H(r,\theta)\,e^{i(kz-\omega t)}$ (S. 5)
where $z$ axis is parallel ($||^{n}$) to the cylinders long axis. More compactly, we can write
$\nabla^{2}_{r,\theta}E+u^{2}E=0\quad\&\quad\nabla^{2}_{r,\theta}H+u^{2}H=0$ (S. 6)
where $\nabla^{2}_{r,\theta}$ is the two-dimensional Laplacian in polar coordinates and $u^{2}=\Big(\frac{\omega^{2}}{c^{2}}\mu(\omega)\varepsilon(\omega)-k^{2}\Big)$. A general solution of the magnetic field in Eq. (S. 6) ($r<a$) can be obtained in cartesian coordinates for one cylinder placed at the origin as shown in Fig. (1) in the main text,
$H^{\text{in}}_{z}=\sum_{n}B_{n}J_{n}(ur)e^{in\theta}$ (S. 7a)
$H^{\text{in}}_{x}=\sum_{n}X_{n}J_{n}(ur)e^{in\theta}$ (S. 7b)
$H^{\text{in}}_{y}=\sum_{n}Y_{n}J_{n}(ur)e^{in\theta}$ (S. 7c)
Our main goal is to represent the radial and tangential components of magnetic field, $H^{\text{in}}_{r}$ and $H^{\text{in}}_{\theta}$ in terms of the cartesian components $H^{\text{in}}_{x}$ and $H^{\text{in}}_{y}$ in Eqs. (S. 7b) and (S. 7c). Following standard notation, we can easily write
$H^{\text{in}}_{r}=H^{\text{in}}_{x}\cos(\theta)+H^{\text{in}}_{y}\sin(\theta)$ (S. 8)
$H^{\text{in}}_{\theta}=-H^{\text{in}}_{x}\sin(\theta)+H^{\text{in}}_{y}\cos(\theta)$ (S. 9)
Using Eqs. (S. 7b) and (S. 7c), we rewrite Eqs. (S. 8) and (S. 9) as,
$H^{\text{in}}_{r}=\sum_{n}\Big(X_{n}\cos(\theta)+Y_{n}\sin(\theta)\Big)J_{n}(ur)e^{in\theta}$ (S. 10)
$H^{\text{in}}_{\theta}=\sum_{n}\Big(Y_{n}\cos(\theta)-Y_{n}\sin(\theta)\Big)J_{n}(ur)e^{in\theta}$ (S. 11)
With the help of $\cos(\theta)=\frac{e^{i\theta}+e^{-i\theta}}{2}$ and $\sin(\theta)=\frac{e^{i\theta}-e^{-i\theta}}{2i}$, the radial component of the magnetic field in Eq. (S. 10) can be simplified as,
$H^{\text{in}}_{r}=\sum_{n}\Bigg[\Big(\frac{X_{n}e^{i(n+1)\theta}-iY_{n}e^{i(n+1)\theta}}{2}\Big)J_{n}(ur)+\Big(\frac{X_{n}e^{i(n-1)\theta}+iY_{n}e^{i(n-1)\theta}}{2}\Big)J_{n}(ur)\Bigg]$ (S. 12)
Replace $n$ by n1n-1 in the first term and $n$ by $n+1$ in the second term of Eq. (S. 12), yields
$H^{\text{in}}_{r}=\sum_{n}\Big[W_{n}J_{n-1}(ur)+Z_{n}J_{n+1}(ur)\Big]e^{in\theta}$ (S. 13)
where $W_{n}=\frac{1}{2}\Big(X_{n-1}-iY_{n-1}\Big)$ and $Z_{n}=\frac{1}{2}\Big(X_{n+1}+iY_{n+1}\Big)$. Similarly we obtain a simplified expression for tangential component $H^{\text{in}}_{\theta}$ in Eq. (S. 11) as,
$H^{\text{in}}_{\theta}=i\sum_{n}\Big[W_{n}J_{n-1}(ur)-Z_{n}J_{n+1}(ur)\Big]e^{in\theta}$ (S. 14)
As $B_{n}$, $W_{n}$ and $Z_{n}$ are not independent parameters, we further connect them using the Maxwell equation $\nabla\cdot H=0$, which in cylindrical coordinate can be expressed as,
$\frac{1}{r}\frac{\partial}{\partial r}(rH_{r})+\frac{1}{r}\frac{\partial H_{\theta}}{\partial\theta}+ikH_{z}=0$ (S. 15)
After substituting the expressions for $H_{r}$, $H_{\theta}$ and $H_{z}$ into Eq. (S. 15), we obtain
$\frac{1}{ur}\Big[W_{n}J_{n-1}+Z_{n}J_{n+1}-nW_{n}J_{n-1}+nW_{n}J_{n+1}\Big]+W_{n}J^{\prime}_{n-1}+Z_{n}J^{\prime}_{n+1}+\frac{ik}{u}B_{n}J_{n}=0$ (S. 16)
Now we can use the recurrence relation of ordinary Bessel functions $J_{n}$, $J^{\prime}_{n+1}(x)=\frac{1}{2}[J_{n-1}(x)-J_{n+1}(x)]$ , $\forall x$, that yields a compact relation between the coefficients $B_{n}$, $W_{n}$ and $Z_{n}$ as,
$W_{n}+Z_{n}+\frac{ik}{u}B_{n}=0$ (S. 17)
We know the $z$ component of electric field can be presented as,
$E_{z}=\sum_{n}A_{n}J_{n}(ur)e^{in\theta}$ (S. 18)
which is connected to the radial and tangential components of magnetic field $H_{r}$ and $H_{\theta}$ by another Maxwell equation
$(\nabla\times H)_{z}=-\frac{i\omega\varepsilon}{c}E_{z}$ (S. 19)
Again in cylindrical coordinate, Eq. (S. 19) can be written as
$\frac{1}{r}\Big(H_{\theta}-\frac{\partial H_{r}}{\partial\theta}\Big)+\frac{\partial H_{\theta}}{\partial r}=-\frac{i\omega\varepsilon}{c}E_{z}$ (S. 20)
Similarly by substituting the expressions of $H_{r}$, $H_{\theta}$ and $E_{z}$ into Eq. (S. 20), we extract a relation between the coefficients $A_{n}$, $W_{n}$ and $Z_{n}$ after using the recurrence relation of ordinary Bessel functions,
$W_{n}+Z_{n}=\frac{\omega\varepsilon}{cu}A_{n}$ (S. 21)
Eqs. (S. 17) and (S. 21) determine the coefficients of $H_{r}$ and $H_{\theta}$ in terms of $A_{n}$ and $B_{n}$, the coefficients of the $z$ component of electric and magnetic field. After solving these two equations, we obtain
$W_{n}=\frac{1}{2}\Big(\frac{\omega\varepsilon}{cu}A_{n}+\frac{ik}{u}B_{n}\Big)\quad\&\quad Z_{n}=\frac{1}{2}\Big(\frac{\omega\varepsilon}{cu}A_{n}-\frac{ik}{u}B_{n}\Big)$ (S. 22)
These expressions of $W_{n}$ and $Z_{n}$ in Eq. (S. 22) lead to a set of normalized solutions for some of the components of $E$ and $H$ inside the cylindrical region from Eq. (S. 6) as,
$E^{\text{in}}_{z}=\sum_{n=-\infty}^{\infty}A_{n}\frac{J_{n}(u_{1}r)}{J_{n}(u_{1}a)}e^{in\theta}$ (S. 23a)
$H^{\text{in}}_{z}=\sum_{n=-\infty}^{\infty}B_{n}\frac{J_{n}(u_{1}r)}{J_{n}(u_{1}a)}e^{in\theta}$ (S. 23b)
$H^{\text{in}}_{r}=\sum_{n=-\infty}^{\infty}\frac{1}{u_{1}}\Bigg(A_{n}\frac{n\omega\varepsilon_{1}}{cu_{1}r}\frac{J_{n}(u_{1}r)}{J_{n}(u_{1}a)}+B_{n}ik\frac{J^{\prime}_{n}(u_{1}r)}{J_{n}(u_{1}a)}\Bigg)e^{in\theta}$ (S. 23c)
$H^{\text{in}}_{\theta}=\sum_{n=-\infty}^{\infty}\frac{i}{u_{1}}\Bigg(A_{n}\frac{\omega\varepsilon_{1}}{c}\frac{J^{\prime}_{n}(u_{1}r)}{J_{n}(u_{1}a)}+B_{n}\frac{ikn}{u_{1}r}\frac{J_{n}(u_{1}r)}{J_{n}(u_{1}a)}\Bigg)e^{in\theta},$ (S. 23d)
where $u^{2}_{1}(\omega)=\Big(\frac{\omega^{2}}{c^{2}}\mu_{1}(\omega)\varepsilon_{1}(\omega)-k^{2}\Big)$, ‘1’ denotes the coordinates of cylinder 1 and $J_{n}$ is the ordinary Bessel function of first kind of order $n$ in standard mathematical notation. Similarly, the corresponding solutions outside of cylinder 1 are,
$E^{\text{ext}}_{z}=\sum_{n=-\infty}^{\infty}\Bigg(C_{n}\frac{H_{n}(u_{3}r)}{H_{n}(u_{3}a)}+F_{n}\frac{J_{n}(u_{3}r)}{J_{n}(u_{3}a)}\Bigg)e^{in\theta}$ (S. 24a)
$H^{\text{ext}}_{z}=\sum_{n=-\infty}^{\infty}\Bigg(D_{n}\frac{H_{n}(u_{3}r)}{H_{n}(u_{3}a)}+G_{n}\frac{J_{n}(u_{3}r)}{J_{n}(u_{3}a)}\Bigg)e^{in\theta}$ (S. 24b)
$H^{\text{ext}}_{r}=\sum_{n=-\infty}^{\infty}\frac{1}{u_{3}}\Bigg(C_{n}\frac{n\omega\varepsilon_{3}}{cu_{3}r}\frac{H_{n}(u_{3}r)}{H_{n}(u_{3}a)}+D_{n}ik\frac{H^{\prime}_{n}(u_{3}r)}{H_{n}(u_{3}a)}+F_{n}\frac{n\omega\varepsilon_{3}}{cu_{3}r}\frac{J_{n}(u_{3}r)}{J_{n}(u_{3}a)}+G_{n}ik\frac{J^{\prime}_{n}(u_{3}r)}{J_{n}(u_{3}a)}\Bigg)e^{in\theta}$ (S. 24c)
$H^{\text{ext}}_{\theta}=\sum_{n=-\infty}^{\infty}\frac{i}{u_{3}}\Bigg(C_{n}\frac{\omega\varepsilon_{3}}{c}\frac{H^{\prime}_{n}(u_{3}r)}{H_{n}(u_{3}a)}+D_{n}\frac{ikn}{u_{3}r}\frac{H_{n}(u_{3}r)}{H_{n}(u_{3}a)}+F_{n}\frac{\omega\varepsilon_{3}}{c}\frac{J^{\prime}_{n}(u_{3}r)}{J_{n}(u_{3}a)}+G_{n}\frac{ikn}{u_{3}r}\frac{J_{n}(u_{3}r)}{J_{n}(u_{3}a)}\Bigg)e^{in\theta}$ (S. 24d)
where $u^{2}_{3}(\omega)=\Big(\frac{\omega^{2}}{c^{2}}\mu_{3}(\omega)\varepsilon_{3}(\omega)-k^{2}\Big)$ and $H_{n}$ is the Hankel function of first kind of order $n$ in standard mathematical notation. Here ‘3’ denotes the background media where the second terms of Eqs. (S. 24a) and (S. 24b) are required to satisfy the boundary conditions at the interfaces due to the presence of cylinder 2. All coefficients $A_{n}$, $B_{n}$, $C_{n}$, $D_{n}$, $F_{n}$ and $G_{n}$ are connected via boundary conditions at the edge of the cylinders $r=a$.
It may be useful to explore in some more detail the physical content of these expansions. The reason why only the ordinary Bessel functions $J_{n}$ occur inside the cylindrical region, is that the fields are required to be finite at the center $r=0$. In the outside region, the coefficients $C_{n}$ and $D_{n}$ describe the outward moving wave, while the $F_{n}$ and $G_{n}$ coefficients $F_{n}$ capture the remaining fields. Here, $F_{n}$ and $G_{n}$ correspond to the external response to the incident wave. In all cases, the Bessel functions are solutions to the governing field equation, while the occurrence of two integration constants reflects the second order property of this equation.
The coefficients $F_{n}$ and $G_{n}$ represent the incident wave, whereas the coefficients $C_{n}$ and $D_{n}$ describe the reflected wave, which are the external responses to the incident wave. To determine these coefficients, it is necessary to ensure the continuity of $E_{z}$, $H_{z}$, $H_{\theta}$, and $\mu H_{r}$. These boundary conditions yield the following expressions,
$A_{n}=C_{n}+F_{n}\quad\&\quad B_{n}=D_{n}+G_{n}$ (S. 25)
$-\frac{nk}{u^{2}_{1}a}B_{n}+\frac{i\omega\varepsilon_{1}}{cu_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{1}a)}A_{n}=\frac{i\omega\varepsilon_{3}}{cu_{3}}\frac{H^{\prime}_{n}(u_{3}a)}{J_{n}(u_{3}a)}C_{n}+\frac{i\omega\varepsilon_{3}}{cu_{3}}\frac{J^{\prime}_{n}(u_{3}a)}{J_{n}(u_{3}a)}F_{n}-\frac{nk}{u^{2}_{3}a}D_{n}-\frac{nk}{u^{2}_{3}a}G_{n}$ (S. 26)
$\mu_{1}\Bigg[\frac{n\omega\varepsilon_{1}}{cu^{2}_{1}a}A_{n}+\frac{ik}{u_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{1}a)}B_{n}\Bigg]=\mu_{3}\Bigg[\frac{n\omega\varepsilon_{3}}{cu^{2}_{3}a}C_{n}+\frac{ik}{u_{3}}\frac{H^{\prime}_{n}(u_{3}a)}{H_{n}(u_{3}a)}D_{n}+\frac{n\omega\varepsilon_{3}}{cu^{2}_{3}a}F_{n}+$ (S. 27)
$\frac{ik}{u_{3}}\frac{J^{\prime}_{n}(u_{3}a)}{J_{n}(u_{3}a)}G_{n}\Bigg]$
where primed Bessel functions are derivative with respect to the radial coordinates. Now we can easily eliminate the $A_{n}$ and $B_{n}$ from Eqs. (S. 26) and (S. 27) using Eq. (S. 25), then we express $C_{n}$ and $D_{n}$ in terms of $F_{n}$ and $G_{n}$ in a matrix form as,
$\begin{pmatrix}C_{n}\\ D_{n}\end{pmatrix}=\begin{pmatrix}Q^{11}_{n}&Q^{12}_{n}\\ Q^{21}_{n}&Q^{22}_{n}\end{pmatrix}\begin{pmatrix}F_{n}\\ G_{n}\end{pmatrix}$ (S. 28)
where the matrix elements are
$Q^{11}_{n}=\Bigg(\frac{\alpha\delta^{\prime}-\alpha^{\prime}\delta}{\alpha\beta^{\prime}-\alpha^{\prime}\beta}\Bigg)\quad\&\quad Q^{12}_{n}=\Bigg(\frac{\alpha^{\prime}\gamma-\alpha\gamma^{\prime}}{\alpha\beta^{\prime}-\alpha^{\prime}\beta}\Bigg),$ (S. 29)
$Q^{21}_{n}=\Bigg(\frac{\beta\delta^{\prime}-\beta^{\prime}\delta}{\alpha\beta^{\prime}-\alpha^{\prime}\beta}\Bigg)\quad\&\quad Q^{22}_{n}=\Bigg(\frac{\beta^{\prime}\gamma-\beta\gamma^{\prime}}{\alpha\beta^{\prime}-\alpha^{\prime}\beta}\Bigg)$
and
$\gamma=-\alpha=\frac{kn}{a}\Big(\frac{1}{u^{2}_{1}}-\frac{1}{u^{2}_{3}}\Big),\quad\quad\beta^{\prime}=-\delta^{\prime}=\frac{n\omega}{ca}\Big(\frac{\varepsilon_{3}\mu_{3}}{u^{2}_{3}}-\frac{\varepsilon_{1}\mu_{1}}{u^{2}_{1}}\Big)$ (S. 30)
$\alpha^{\prime}=ik\Big(\frac{\mu_{1}}{u_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{1}a)}-\frac{\mu_{3}}{u_{3}}\frac{H^{\prime}_{n}(u_{3}a)}{H_{n}(u_{3}a)}\Big),\quad\gamma^{\prime}=ik\Big(\frac{\mu_{3}}{u_{3}}\frac{J^{\prime}_{n}(u_{3}a)}{J_{n}(u_{3}a)}-\frac{\mu_{1}}{u_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{1}a)}\Big)$
$\beta=\frac{i\omega}{c}\Big(\frac{\varepsilon_{3}}{u_{3}}\frac{H^{\prime}_{n}(u_{3}a)}{H_{n}(u_{3}a)}-\frac{\varepsilon_{1}}{u_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{1}a)}\Big),\quad\delta=\frac{i\omega}{c}\Big(\frac{\varepsilon_{1}}{u_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{3}a)}-\frac{\varepsilon_{3}}{u_{3}}\frac{J^{\prime}_{n}(u_{3}a)}{J_{n}(u_{3}a)}\Big)$
The above expressions completely determine the response of a single cylinder with radius $a$ to an arbitrary incident electromagnetic wave. For nonmagnetic medium, the matrix in Eq. (S. 28) agrees with previously found results by Mitchell et al.23.
At this point we consider the case when both the cylinders are present in the background media. It is noted that the arbitrary incident waves with coefficients $F_{n}$ and $G_{n}$ in the first cylinder are linked to a reflected wave coming from the second cylinder. The $z$ components of the reflected fields from the second cylinder can be written as,
$E_{z}=\sum_{n=-\infty}^{\infty}C^{\prime}_{n}\frac{H_{n}(u_{3}r_{2})}{H_{n}(u_{3}a)}e^{in\theta_{2}}$ (S. 31)
$H_{z}=\sum_{n=-\infty}^{\infty}D^{\prime}_{n}\frac{H_{n}(u_{3}r_{2})}{H_{n}(u_{3}a)}e^{in\theta_{2}}$
To correlate between the incident waves onto the first cylinder and the reflected waves from the second cylinder we need to express $F_{n}$ and $G_{n}$ in terms of $C^{\prime}_{n}$ and $D^{\prime}_{n}$. The general strategy to do this is as follows: first we need to express the reflected fields of the second cylinder in the coordinates of the first cylinder ($r,\theta,z$) using Graf’s addition theorem 40 for Hankel function,
$H_{n}(u_{3}r_{2})\,e^{in\theta_{2}}=\sum_{m=-\infty}^{\infty}H_{m+n}(u_{3}R)\,J_{m}(u_{3}r)\,e^{im\theta}$ (S. 32)
Hence the magnetic field in Eq. (15) can be written as
$H_{z}=\sum_{m=-\infty}^{\infty}\frac{J_{m}(u_{3}r)}{J_{m}(u_{3}a)}J_{m}(u_{3}a)\sum_{n=-\infty}^{\infty}D^{\prime}_{n}\frac{H_{m+n}(u_{3}R)}{H_{n}(u_{3}a)}e^{im\theta}$ (S. 33)
We can now easily identify the coefficients of $e^{im\theta}$ with the coefficient $G_{m}$ by comparing Eqs. (S. 24b) and (S. 33) that yields
$G_{m}=J_{m}(u_{3}a)\sum_{n=-\infty}^{\infty}D^{\prime}_{n}\frac{H_{m+n}(u_{3}R)}{H_{n}(u_{3}a)}$ (S. 34)
Similarly we obtain a further equation for the coefficient $F_{m}$ of the $z$ component of electric field as,
$F_{m}=J_{m}(u_{3}a)\sum_{n=-\infty}^{\infty}C^{\prime}_{n}\frac{H_{m+n}(u_{3}R)}{H_{n}(u_{3}a)}$ (S. 35)
Using these expressions, Eq. (S. 28) is written as,
$\begin{pmatrix}C_{m}\\ D_{m}\end{pmatrix}=\underbrace{\sum_{n}Q_{mn}\begin{pmatrix}Q^{11}_{m}&Q^{12}_{m}\\ Q^{21}_{m}&Q^{22}_{m}\end{pmatrix}}_{M_{mn}}\begin{pmatrix}C^{\prime}_{n}\\ D^{\prime}_{n}\end{pmatrix}$ (S. 36)
where
$Q_{mn}=J_{m}(u_{3}a)\frac{H_{m+n}(u_{3}R)}{H_{n}(u_{3}a)}$ (S. 37)
Now let’s consider $\widetilde{X}$ as an infinite column matrix where the matrix elements are   $\cdots C_{m},\,D_{m}\cdots$ and a infinite matrix $\widetilde{M}$ with corresponding matrix elements are $M_{mn}$ and another infinite column matrix $\widetilde{X}^{\prime}$ whose elements are   $\cdots C^{\prime}_{n},\,D^{\prime}_{n}\cdots$. A compact representation of Eq. (S. 36) is,
$\widetilde{X}=\widetilde{M}\widetilde{X}^{\prime}$ (S. 38)
A similar type of expression exists when the interaction is viewed from the second cylinder,
$\widetilde{X}^{\prime}=\widetilde{N}\widetilde{X}$ (S. 39)
where $\widetilde{N}$ can be deduced from $\widetilde{M}$ upon replacing $u_{1}$ by $u_{2}$, $u^{2}_{2}(\omega)=\Big(\frac{\omega^{2}}{c^{2}}\mu_{2}(\omega)\varepsilon_{2}(\omega)-k^{2}\Big)$. From combing the above equations we are able to obtain the dispersion relation of the surface modes:
$\mathscr{D}(\omega)\equiv\mathrm{Det}(\mathbb{I}-\widetilde{\Omega})=0;\quad\widetilde{\Omega}=\widetilde{M}\widetilde{N}$ (S. 40)
Hence the free energy of interaction per unit length (for a cylinder with length $L$) in terms of the dispersion relation is given by
$G(a,R,T)=\frac{k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\,\ln\mathscr{D}(i\xi_{p})$ (S. 41)
where the prime on the summation means the term for which $p=0$ must be weighted by $1/2$ and the Matsubara frequency $\xi_{p}=2\pi kTp/\hbar$, $p$ is an integer.
S.2 Dispersion interaction for thin cylinders
Though the expression for the dispersion relation in Eq. (S. 41) defines the full solution of our problem, it is very cumbersome and of no practical advantage in its original form. For more useful and transparent solutions we consider the special case of “thin cylinder approximation” in the limit of $a\ll R$. There exists here a physical lower limit on the magnitude of $a$: it must be so large that classical fluid mechanics is applicable. Usually, a lower limit of about 100 nm should be more than sufficient. One must observe that in the limit of very thin cylindrical structures, e.g. DNA strings, one must use a quantum mechanical formalism.
To reduce the complexity of our solutions, we followed a method developed for finding the non-retarded interaction between cylinders 23. The main contribution comes from the range of axial wave vectors, $k\lesssim 1/R$, which along with the thin cylinder approximation $a/R\ll 1$ leads to a further approximation $|u_{\alpha}a|\ll 1,(\alpha=1,2,3)$ ($|\xi a/c|\ll 1$). In this region, the Bessel functions are expanded to a leading order term for its small arguments $u_{\alpha}a$ and $u_{\alpha}R$ and the approximated matrix elements $M_{mn}$ can be written as,
$M_{mn}\simeq\Big(\frac{a}{R}\Big)^{m+n}\frac{(m+n-1)!}{m!(n-1)!}<\Big(\frac{a}{R}\Big)^{m+n}2^{m+n}$ (S. 42)
This shows that with increasing $m$ and $n$, these matrix elements decrease rapidly. On the other hand when $k>1/R$, using Matsubara frequencies, we find
$u^{2}_{\alpha}(i\xi_{p})=-\Big(\frac{\xi^{2}_{p}}{c^{2}}\mu_{\alpha}(i\xi_{p})\varepsilon_{\alpha}(i\xi_{p})+k^{2}\Big)$ (S. 43)
At this point, it is convenient to introduce a new variable $\nu_{\alpha}=e^{i\pi/2}u_{\alpha}$ leading to a new relation for the modified Bessel function,
$H_{n}(xe^{i\pi/2})=-\frac{2i}{\pi}\,e^{-in\pi/2}K_{n}(x),\quad-\pi<\mathrm{arg}(x)<\pi/2$ (S. 44)
Hence the matrix elements $M_{mn}$ can be connected with the modified Bessel function, using the new variable $\nu_{\alpha}$,
$M_{mn}\propto H_{m+n}(u_{3}R)\propto K_{m+n}(\nu_{3}R)$ (S. 45)
Therefore, in the region $kR\gg 1$ i.e., $|\nu_{3}R|\gg 1$, the matrix elements become
$M_{mn}\simeq\sqrt{\frac{1}{\nu_{3}R}}\,e^{-\nu_{3}R}$ (S. 46)
The above asymptotic expression, shows that in this limit the matrix elements are finite and they are independent of $m$ and $n$.
Thus, we obtain finite contributions in the long and short separation limits. From Eqs. (S. 42) and (S. 46), it is very clear that the dominant effect can only be found in lower values of $m$ and $n$; i.e., $m,n=0,\pm 1$. The leading term which we require is then obtained by expanding the logarithmic term in Eq. (S. 41) as,
[…]
$\simeq-[M_{00}\,N_{00}+2M_{11}\,N_{11}+2M_{10}\,N_{01}+2M_{01}\,N_{10}+2M_{1-1}\,N_{-11}]$
Using Eqs. (S. 36) and (S. 39), the matrix elements can be written down as,
$-[M_{00}\,N_{00}+2M_{11}\,N_{11}+2M_{10}\,N_{01}+2M_{01}\,N_{10}+2M_{1-1}\,N_{-11}]=-Q_{00}S_{00}(Q^{11}_{0}S^{11}_{0}+Q^{22}_{0}S^{22}_{0})$ (S. 48)
$-2Q_{01}S_{10}(Q^{11}_{0}S^{11}_{1}+Q^{22}_{0}S^{22}_{1})-2Q_{10}S_{01}(Q^{11}_{1}S^{11}_{0}+Q^{22}_{1}S^{22}_{0})-2(Q_{11}S_{11}+Q_{1,-1}S_{-1,1})$
$(Q^{11}_{1}S^{11}_{1}+Q^{12}_{1}S^{21}_{1}+Q^{21}_{1}S^{12}_{1}+Q^{22}_{1}S^{22}_{1})$
where the functions $S$ are determined from the functions $Q$ by replacing $\varepsilon_{1}$ and $\mu_{1}$ by $\varepsilon_{2}$ and $\mu_{2}$ respectively. It is worth mentioning that when the order of the ordinary Bessel function and Hankel function is not equal zero, then they can be approximated as,
$\frac{J^{\prime}_{n}(x)}{J_{n}(x)}\simeq\frac{n}{x}\quad\quad\frac{H^{\prime}_{n}(x)}{H_{n}(x)}\simeq-\frac{n}{x}$ (S. 49)
Using limiting forms of Bessel functions and Hankel functions, we list the values of matrix elements in Eq. (S. 36) in the following way:
$Q^{11}_{n}$ $=-\Bigg[\frac{\varepsilon_{1}(\mu_{1}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{1})}{u^{2}_{1}}-\frac{2\varepsilon_{3}\mu_{3}}{u^{2}_{3}}\Bigg]\frac{u^{2}_{1}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})},\quad n\neq 0$ (S. 50)
$=\Bigg[\frac{\varepsilon_{3}-\varepsilon_{1}}{2\varepsilon_{3}}\Bigg]a^{2}u^{2}_{3}\ln(u_{3}a),\quad n=0$
$Q^{12}_{n}$ $=\Bigg[\frac{ikc\mu_{3}u^{2}_{1}}{\omega(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})}\Bigg]\Bigg[\frac{1}{u^{2}_{1}}-\frac{1}{u^{2}_{3}}\Bigg],\quad n\neq 0$ (S. 51)
$=0,\quad n=0$
$Q^{21}_{n}$ $=\frac{i\omega}{kc}\Bigg[\frac{\varepsilon_{1}\mu_{1}}{u^{2}_{1}}-\frac{\varepsilon_{3}\mu_{3}}{u^{2}_{3}}\Bigg]\frac{2\varepsilon_{3}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})}\,u^{2}_{1},\quad n\neq 0$ (S. 52)
$=0,\quad n=0$
$Q^{22}_{n}$ $=-\Bigg[\frac{\varepsilon_{3}(\mu_{1}+\mu_{3})+\varepsilon_{1}(\mu_{1}-\mu_{3})}{u^{2}_{1}}-\frac{2\varepsilon_{3}\mu_{3}}{u^{2}_{3}}\Bigg]\frac{u^{2}_{1}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})},\quad n\neq 0$ (S. 53)
$=\Bigg[\frac{\mu_{3}-\mu_{1}}{2\mu_{3}}\Bigg]a^{2}u^{2}_{3}\ln(u_{3}a),\quad n=0$
when $x\to 0$, $K_{0}(x)\simeq-\ln(x)$. Therefore the elements corresponding to Eq. (S. 37) are expressed as,
$\begin{split}Q_{00}=-\frac{i\pi}{2}\frac{H_{0}(u_{3}R)}{\ln(u_{3}a)},\quad\&\quad Q_{01}=ia\pi\Big(\frac{u_{3}}{2}\Big)H_{0}(u_{3}R)\\ Q_{10}=-a\Big(\frac{i\pi u_{3}}{4}\Big)\frac{H_{0}(u_{3}R)}{\ln(u_{3}a)},\quad\&\quad Q_{11}=a^{2}\Big(\frac{i\pi u^{2}_{3}}{4}\Big)H_{2}(u_{3}R)\\ Q_{1,-1}=-ia^{2}\pi\Big(\frac{u_{3}}{2}\Big)^{2}H_{0}(u_{3}R)\end{split}$ (S. 54)
and
$Q_{11}=-Q_{-1,-1};\quad Q_{-1,1}=Q_{1,-1}$ (S. 55)
$Q_{01}=Q_{0,-1};\quad\quad Q_{10}=Q_{-1,0}$
Furthermore, using Eqs. (S. 48) and (S. 50)-(S. 55), Eq. (S. 47) can be written as,
$\ln\mathscr{D}(i\xi_{p})=f_{0}K^{2}_{0}(\nu_{3}R)+f_{1}K^{2}_{1}(\nu_{3}R)+f_{2}\Big(K^{2}_{0}(\nu_{3}R)+K^{2}_{2}(\nu_{3}R)\Big)$ (S. 56)
where $K_{0},K_{1}$ and $K_{2}$ are modified Bessel functions of zeroth, first and second orders respectively, and the coefficients are as follows:
$f_{0}=-\frac{a^{4}}{4}\Bigg[\frac{(\varepsilon_{3}-\varepsilon_{1})(\varepsilon_{3}-\varepsilon_{2})}{\varepsilon^{2}_{3}}+\frac{(\mu_{3}-\mu_{1})(\mu_{3}-\mu_{2})}{\mu^{2}_{3}}\Bigg]\nu^{4}_{3}$ (S. 57)
$f_{1}=-a^{4}\Bigg\{\frac{\nu^{2}_{1}\nu^{4}_{3}}{(\varepsilon_{3}+\varepsilon_{1})(\mu_{3}+\mu_{1})}\Bigg[\frac{(\varepsilon_{3}-\varepsilon_{2})}{\varepsilon_{3}}\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{1}(\mu_{1}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{1})}{2\nu^{2}_{1}}\Bigg)+\frac{(\mu_{3}-\mu_{2})}{\mu_{3}}$ (S. 58)
$\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{1}+\mu_{3})+\varepsilon_{1}(\mu_{1}-\mu_{3})}{2\nu^{2}_{1}}\Bigg)\Bigg]+\frac{\nu^{2}_{2}\nu^{4}_{3}}{(\varepsilon_{3}+\varepsilon_{2})(\mu_{3}+\mu_{2})}\Bigg[\frac{(\varepsilon_{3}-\varepsilon_{1})}{\varepsilon_{3}}$
$\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{2}(\mu_{2}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{2})}{2\nu^{2}_{2}}\Bigg)+\frac{(\mu_{3}-\mu_{1})}{\mu_{3}}\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{2}+\mu_{3})+\varepsilon_{2}(\mu_{2}-\mu_{3})}{2\nu^{2}_{2}}\Bigg)\Bigg]\Bigg\}$
$f_{2}=-\frac{a^{4}\nu^{2}_{1}\nu^{2}_{2}\nu^{4}_{3}}{(\varepsilon_{3}+\varepsilon_{1})(\varepsilon_{3}+\varepsilon_{2})(\mu_{3}+\mu_{1})(\mu_{3}+\mu_{2})}\Bigg[\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{1}(\mu_{1}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{1})}{2\nu^{2}_{1}}\Bigg)$ (S. 59)
$\times\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{2}(\mu_{2}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{2})}{2\nu^{2}_{2}}\Bigg)+\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{1}}\Bigg)\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{2}\mu_{2}}{\nu^{2}_{2}}\Bigg)$
$+\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{2}}\Bigg)\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{1}\mu_{1}}{\nu^{2}_{1}}\Bigg)+\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{1}+\mu_{3})+\varepsilon_{1}(\mu_{1}-\mu_{3})}{2\nu^{2}_{1}}\Bigg)\times$
$\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{2}+\mu_{3})+\varepsilon_{2}(\mu_{2}-\mu_{3})}{2\nu^{2}_{2}}\Bigg)\Bigg]$
S.3 Excited state interaction derivation
In the case of two identical thin cylindrical nanoparticles, if a long-lived antisymmetric state exists it results in a long-ranged excited state interaction. In general, this would formally be given by 26,
$G^{\text{res}}(a,R,T)=\frac{2k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\ln[{\mathrm{Det}(\mathbb{I}-M)}]=-\frac{2k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\,\mathrm{Tr}[M_{mn}]$ (S. 60)
We are only interested in the first few values of $m,n=0,\pm 1$ in $M_{mn}$, as matrix elements decrease rapidly with increasing $m$ and $n$ (the follow an exponential decay in thin cylinder approximation and $kR\gg 1$). The trace of the matrix elements $M_{mn}$ can be written from its definition as,
$\mathrm{Tr}[M_{mn}]=\sum_{m,n=-1}^{1}Q_{mn}\big(Q^{11}_{m}+Q^{22}_{m}\big)$ (S. 61)
Ignoring the imaginary terms, we can write down all the matrix elements explicitly,
$\mathrm{Tr}[M_{mn}]=2Q^{11}_{-1}Q_{1,-1}+2Q^{22}_{-1}Q_{1,-1}+Q_{00}\,Q^{11}_{0}+Q_{00}\,Q^{22}_{0}$ (S. 62)
A relevant expression for the fully retarded excited-state interaction using Eqs. (S. 50)-(S. 55) and $H_{0}(u_{3}R)=-\frac{2i}{\pi}K_{0}(\nu_{3}R),\,\,-\pi<\mathrm{arg}(u_{3})<\pi/2$, $\nu^{2}_{\alpha}=-u^{2}_{\alpha},\quad\alpha=1,2,3$ is
$G^{\text{res}}(a,R,T)$ $=-\frac{2k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\,(t_{0}+t_{1}+t_{2}+t_{3})\,K_{0}(\nu_{3}R)$ (S. 63)
where
$t_{0}=\Bigg[\frac{\varepsilon_{3}-\mu_{1}}{2\,\varepsilon_{3}}\Bigg]a^{2}\nu^{2}_{3};\quad\quad t_{1}=\Bigg[\frac{\mu_{3}-\varepsilon_{1}}{2\,\mu_{3}}\Bigg]a^{2}\nu^{2}_{3}$ (S. 64)
$t_{3}=\Bigg[-\frac{\varepsilon_{1}(\mu_{1}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{1})}{\nu^{2}_{1}}+\frac{2\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}\Bigg]\frac{a^{2}\nu^{2}_{1}\nu^{2}_{3}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})}$
$t_{4}=\Bigg[-\frac{\varepsilon_{3}(\mu_{1}+\mu_{3})+\varepsilon_{1}(\mu_{1}-\mu_{3})}{\nu^{2}_{1}}+\frac{2\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}\Bigg]\frac{a^{2}\nu^{2}_{1}\nu^{2}_{3}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})}$
It is well-known that at large separation, the maximum contribution for the interaction energy comes from the classical zero frequency (entropic) term whose long-range interaction is
$G^{\text{res}}_{p=0}(a,R,T)=-\frac{k_{B}T}{\pi}\int_{0}^{\infty}dk\,(t_{0}+t_{1}+t_{2}+t_{3})\,K_{0}(\nu_{3}R)$ (S. 65)
S.4 Preliminary considerations for the impact of magnetic fluid
In view of the complexity of the formalism, we first note some relative magnitudes of central quantities as part of the theory. Consider first the dispersion relation: adopting a Lorentz form for the ferrofluid magnetic susceptibility $\chi_{ms}$ with a characteristic frequency $\omega_{0m}$, we have 25
$\mu(i\xi)=1+\frac{4\pi\chi_{ms}}{1+\xi^{2}/\omega_{0m}^{2}},$ (S. 66)
with $\xi=-i\omega$ being the imaginary frequency. Here $\omega_{0m}$ is relatively low, typically $\omega_{0m}=3\times 10^{10}$ rad/s, in contrast to the characteristic dielectric $\omega_{0e}$ which is typically $\omega_{0e}=3\times 10^{16}$ rad/s. Hence, only zero frequency contribution is influenced strongly by the magnetic effects.
At large separations, non-zero frequency contributions to the dispersion interaction are screened out by retardation. Thus, the magnetic part of the dispersion forces is an essential part of the long-range/high-temperature zero-frequency contribution. A characteristic property of both the non-retarded and zero-frequency entropic interaction between cylinders is its scaling $R^{-5}$ dependence32, 23, 41, 42, 31, 43, 44, 45. Another way of looking at the condition for nonretarded theory is to observe that the distance $R$ between the cylinders must be much smaller than the wavelength $\lambda_{0}=2\pi c/(n\omega_{0})$ corresponding to the proper frequency $\omega_{0}$ for the material ($n$ is the refractive index). In the electric case with $\omega_{0}=\omega_{0e}$ one gets $\lambda_{0e}\sim 50$~nm, while in the magnetic case with $\omega_{0}=\omega_{0m}$ one gets $\lambda_{0m}$ to lie in the centimeter region. Thus in the magnetic case the nonretarded theory is often applicable, at least when the condition $R\ll\lambda_{0}$ is satisfied. Also, in the absence of magnetic fields, effects due to magnetic nonlinearities are usually negligible in Casimir problems. However, a very recent paper discussed how the Casimir force in specific cases can be tuned by magnetic fields 37. The details on how we modeled the magnetic permeability of the magnetic fluid are given in the ”Materials Modeling” section.
S.5 Material modeling
S.5.1 Zinc oxide nanorod
We use zinc oxide (ZnO) 34 as nanorod material as it is a wide-gap semiconductor with many advantages 46, 47, such as: ($i$) availability of large area substrates, ($ii$) high solubility of n-type dopants, (iiiiii) large exciton binding energy, ($iv$) excellent luminescent properties, ($v$) high chemical/thermal stability and radiation hardness, ($vi$) wet chemical synthesis is possible, (viivii) room-temperature ferromagnetism via Mn/Fe-doping, (viiiviii) quantum size effects on exciton and gap energies in nanocrystals, ($ix$) good UV-blocker stability stable, ($x$) low power threshold at room temperature, ($xi$) piezoelectric, (xiixii) stable polar surface, (xiiixiii) outperforming cost-effectiveness, whiteness, and UV-blocking of nanoparticles antibacterial finishing, (xivxiv) provides reinforcement in natural rubber, and ($xv$) it is bio-safe and biocompatible. In addition, high-quality ZnO nanorods and nanowires can be synthesized without widespread defects or dislocations in the crystal structure.
The material properties of ZnO and magnetite (Fe3O4) were investigated within the framework of density functional theory (DFT), employing the projector augmented-wave (PAW) method as implemented in the Vienna Ab initio Simulation Package (VASP)48. To accurately describe the dielectric response at high energies (i.e., extending beyond 3000 Matsubara frequencies), $GW$-type core potentials were employed, an energy cutoff of 600 eV, and the initial valence configurations for the atoms are chosen to be $3s^{2}p^{6}d^{10}4s^{2}$ for Zn, $3s^{2}p^{6}d^{8}$ for Fe, and $2s^{2}p^{4}$ for O. The exchange-correlation interaction for the wide-gap ZnO is modeled with the hybrid functional by Heyd, Scuseria, and Ernzerhof49 with 30% Fock exchange; this yields a band gap energy of 3.34 eV. The hybrid functional generates proper Zn $d$-like orbitals, which is important for describing the valence $sd$-hybridization in ZnO 50. The irreducible Brillouin zone for the 4-atom hexagonal unit cell is sampled by a $12\times 12\times 8$ k-mesh. For the ferrimagnetic Fe3O4 we choose the generalized gradient approximation for solids, developed by Perdew et al.51 with an effective Hubbard correction 52 $U=5.7$ eV on the Fe $d$-like orbitals. A $20\times 20\times 20$ grid of k-points is used for the 14-atom face-centered cubic unit cell. Our calculated dielectric function of Fe3O4 agrees overall very well with the empirical spectrum by Draine and Hensley 53 in the low infrared to visible region. For energies below 1 eV we combine our calculated data with Draine and Hensley’s parameterised spectrum, which then accounts also for the room temperature electrical conductivity.
From the computed electronic structures of the materials, we determine the imaginary part of the dielectric response functions in the long wavelength limit:
$\varepsilon^{\prime\prime}(\omega)=\lim_{{\bf q}\rightarrow 0}\frac{4\pi^{2}e^{2}}{V_{\Omega}q^{2}}\sum_{v,c,{\bf k}}\delta(\epsilon_{c,{\bf k}}-\epsilon_{v,{\bf k}}-\hbar\omega)\langle u_{c,{\bf k}+{\bf e}_{\alpha}q}|u_{v,{\bf k}}\rangle\langle u_{v,{\bf k}}|u_{c,{\bf k}+{\bf e}_{\alpha}q}\rangle$ (S. 67)
$+\sum_{j}\frac{S_{j}\,\omega_{\rm TO,j}^{2}\Gamma_{j}\omega}{(\omega_{{\rm TO},j}^{2}-\omega^{2})^{2}+(\Gamma_{j}\omega)^{2}}+\frac{\Gamma\,\omega_{\rm pl}^{2}}{\omega(\omega^{2}+\Gamma^{2})}$
where the first term describes the electronic response. Here, $V_{\Omega}$ is the unit-cell volume referenced to the three Cartesian directions ${\bf e}_{\alpha}$. The eigenfunctions are represented by the cell periodic part $u_{v/c}$ of the valence ($v$) or conduction ($c$) state with corresponding energy $\epsilon_{v/c,{\bf k}}$. The second term in Eq. (S. 67) describes the ionic response, where the vibrations screen the carriers by an electric field. Here, $\omega_{TO}$ is the frequency of the transverse optical mode, $\Gamma_{j}$ is the damping, and $S_{j}$ is the oscillator strength of the $j$th mode in its vibration direction. From the density functional perturbation theory we compute the Hessian matrix of the ionic displacements to model the optical phonons. The third term in Eq. (S. 67) describes the Drude screening with plasma frequency $\omega_{\rm pl}$ arising from nearly free carriers in the partially occupied bands. The dielectric functions are modeled for the imaginary frequencies and presented in Fig. (2) in the main text.
S.5.2 Polystyrene nanorod and dielectric function of Toluene
The nanoparticles we consider have large enough dimensions to approximate their effective dielectric functions with those of bulk materials. We use available experimental data from the literature for polystyrene (PS) nanorod33 and this reference is also used to provide the dielectric function for Toluene.
S.6 Thermal stability
In this section we have performed stability analysis with respect to the thermal fluctuations, following the detailed procedure outlined in Ref.21.
Figure S.1: Colors online: Thermal stability analysis in the attractive region of total force for diameter $D=17$ nm and $\phi=5\%$.
In this situation, the trapping width is 500 nm and force gradient in the trapping region is $7.92\times 10^{-11}\mu N$. The depth of the corresponding trapping potential is calculated and found to be
$\Delta U\approx 4\times 10^{-23}\mathrm{J}.$
At room temperature ($T=298.15\mathrm{K}$) the thermal energy $k_{B}T\approx 4.11\times 10^{-21}\mathrm{J}$. Comparing the two gives
$\frac{\Delta U}{k_{B}T}\approx 0.01\Rightarrow\Delta U\ll k_{B}T.$
As a consequence, for this specific configuration, we are in a regime where the trapping potential is much reduced compared to the thermal energy and thermal fluctuation can easily surmount the Casimir potential such that the nanoparticle remains unstable. This problem can be enhance using PFA theorem that is beyond our discussion.
S.7 Numerical analysis of excited state interaction
Safari et al.54 used perturbation theory to study dispersion interaction between a pair of chiral molecules in the presence of magnetoelectric media. They found strong magnetic effects. Many biomedical processes occur in background magnetic fluids influencing both ground and excited state interactions. In the long range limit the excited state dispersion interaction, as shown in Fig. (S.2), receives important contributions from the zero frequency term. Notably, excited state interactions between elongated biomolecules can be long-range and impacted by the magnetic permeability.
Figure S.2: Colors online: Contributions to the force per unit length from an antisymmetric excited state between two cylindrical polystyrene nanoparticles with the same radius $a=20$ nm as a function of center-to-center distance $R$ when immersed in a ferrofluid with 5% concentration of magnetite nanoparticles of diameter 17 nm.
S.8 Force calculation with gold-coated magnetite nanoparticles
We study here the interaction forces between ZnO and PS in their ground state when suspended in a magnetic colloidal host formed by gold-coated magnetite nanoparticles dispersed in toluene. The objective of this study is to clarify how the compositional characteristics of magnetic nanoparticles affect the inter-nanorod potential and the resulting trapping behavior.
Figure S.3: Colors online: The CL force per unit length between two cylindrical objects (ZnO and PS respectively) with the same radius $a=20$ nm as a function of center-to-center distance $R$ when immersed in a ferrofluid with 5% concentration of gold-coated (thickness $R_{\rm Au}=2,3,5$ nm) magnetite nanoparticles of diameter 13, 15 and 17 nm.
We find that the trapping between ZnO and PS nanorods (for the considered cases) takes place only for magnetite nanoparticles with core size, $D=17$ nm. This behavior is consistent across all examined gold-layer thicknesses, namely $R_{\rm Au}=2\,\text{nm},3\,\text{nm}$, and 5 nm in Fig. (S.3). We find this trapping to be thermally unstable which we discuss in Sec. (S.6) and can be overcome via PFA theorem.
Data availability: The code and data that support the findings of this study are available in Github.

### v2

@@
 Here, we demonstrate a route to modulate Casmir-Lifshitz interactions between anisotropic nanoparticles with the magnetic fluids.
-By semi-classical quantum electrodynamics, we study ground state dispersion forces for cylindrical dielectric nanorods made of polystyrene (PS), and zinc oxide (ZnO) embedded in toluene-based host media with gold-coated magnetite nanoparticles and also predict magnetic contributions to the non-retarded excited state interaction.
-The variation in magnetic permeability enables tuning between repulsive and attractive interaction and a thermally unstable and measurable magnetic Casimir traps are predicted between a pair of ZnO–PS nanoparticles whose equilibrium position can be modulated over an order of magnitude with a small variation in the size of the magnetite nanoparticle.
+By semi-classical quantum electrodynamics, we study ground state dispersion forces for cylindrical dielectric nanorods made of polystyrene (PS), and zinc oxide (ZnO) embedded in toluene-based host media with gold-coated magnetite nanoparticles and also predict magnetic contributions to the fully retarded excited state interaction.
+The variation in magnetic permeability enables tuning between repulsive and attractive interaction and measurable magnetic Casimir traps are predicted between a pair of ZnO–PS nanoparticles whose equilibrium position can be modulated over an order of magnitude with a small variation in the size of the magnetite nanoparticle.
 This provides an alternative magnetic Casimir-effect pathway to reversibly tune quantum electromagnetic forces at the nanoscale for assembly and enhancement of colloidal stability.
 What if the very surrounding media in space itself could be engineered into a tool for nano-manipulation?
-For instance, can we take the background media as a magnetic fluid, in which electromagnetic fluctuations exert forces that can pull, push or most interestingly trap matters at the nanoscale level?
-Such fluctuation-induced interactions, which are referred to as van der Waals 1, 2 (vdW) and Casimir-Lifshitz (CL) 3, 4 forces together, typically act to bring microscopic bodies into contact, leading to unwanted self-aggregation in nanofabrication and soft-matter systems 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15.
-An important practical issue in nano and micro-devices is to overcome close contact attractions that causes building blocks to cluster together and an ideal solution would be to engineer systems where the interaction turns repulsive at small distances14.
-Beyond simply mitigating stiction, these fluctuations could also be harnessed to produce stable and controllable interaction potentials.
-Fundamental investigations of the influence on CL forces by a magnetic surrounding media have been largely restricted to simple geometries 16, 17, 18.
-Exploiting such a foundation, we have recently expanded the theoretical approach by including the contributions of transverse magnetic (TM) and transverse electric (TE) modes within a model described by the proximity force theorem 19, 20 in such a way that both energy contributions due to TM and TE-modes would partially act as candidates to enable quantum trapping between two quasi-planar surfaces 21.
+For instance, can we take the background media as a magnetic fluid, in which electromagnetic fluctuations exert forces that can pull, push or most interestingly trap matter at the nanoscale level?
+Such fluctuation-induced interactions, which are referred to as van der Waals wang1928problem, Lond (vdW) and Casimir-Lifshitz (CL) casimir1948-1, Dzya forces together, typically act to bring microscopic bodies into contact, leading to unwanted self-aggregation in nanofabrication and soft-matter systems ParsegianNinham1969, dobson2006asymptotics, Munday2009, esteso2015nanolevitation, WoodsRevModPhys.88.045003, zhao2019stable, Woods2020, esteso2022effect, MundayTorque2022, MundayReview2024, Woods2024.
+An important practical issue in nano and micro-devices is to overcome close contact attractions that causes building blocks to cluster together and an experimental solution is to engineer systems where the interaction turns repulsive at small distances Munday2009, esteso2015nanolevitation, WoodsRevModPhys.88.045003, zhao2019stable, Woods2020, esteso2022effect, MundayReview2024, WilliamsonJPCL2025 Beyond simply mitigating stiction, these fluctuations could also be harnessed to produce stable and controllable interaction potentials.
+Fundamental investigations of the influence on CL forces by a magnetic surrounding media have been largely restricted to simple geometries Richmond_1971, klimchitskaya2019impact, VelichkoKlimchitskaya2020.
+Exploiting such a foundation, we have recently expanded the theoretical approach by including the contributions of transverse magnetic (TM) and transverse electric (TE) modes within a model described by the proximity force theorem Derjaguin1934, BLOCKI198153 in such a way that both energy contributions due to TM and TE modes would partially act as candidates to enable quantum trapping between two quasi-planar surfaces pal2025trapping.
 In this letter, we go beyond these simple geometries and study an experimentally feasible setup which comprises of pairs of cylindrical nanorods made of polystyrene (PS) and zinc oxide (ZnO).
 The use of the full cylindrical geometry permits a more realistic description of nanoscale behavior, albeit at the price of considerable added theoretical complexity.
-This stimulates further development of analytical and numerical studies of Lifshitz and vdW interaction between elongated nanoparticles 22, 23, 24, 25, 26, 27.
-One of us previously derived a few asymptotic limits for this kind of a system 25.
+This stimulates further development of analytical and numerical studies of Lifshitz and vdW interaction between elongated nanoparticles Richmond1972, mitchell1973van, Davies1973, NylandBrevik1994, SubhojitPCCP2024, PhysRevE.111.015403.
+One of us previously derived a few asymptotic limits for this kind of a system NylandBrevik1994.
 Here, we examine the magnetic contributions associated with a surrounding ferrofluid for the long-range and high-temperature limits of ground-state dispersion interactions between cylindrical nanoparticles.
@@
 Such a device could enable controlled, magnetic-induced entropic CL trapping.
-Early studies of CL interactions between magnetic fluida 17, 18 were mainly concentrated on the observation of a repulsive force by changing the optical properties of the medium.
+Early studies of CL interactions between magnetic fluida klimchitskaya2019impact, VelichkoKlimchitskaya2020 were mainly concentrated on the observation of a repulsive force by changing the optical properties of the medium.
 By contrast, we have implemented here a way in which this zero-frequency, purely entropically driven Casimir trapping is realized exclusively by tuning the fluid’s magnetic permittivity, an effect not investigated earlier.
@@
 Lastly we demonstrate here that the retardation substantially diminishes the overall interaction strength as well as amplifies the relative weight of magnetic response from the long-range, classical entropic contributions.
-Here, we thus propose a novel experimentally feasible approach to manipulate nanorod pair complexes by suspending them in a magnetically diluted fluid, where the interplay between repulsive non-magnetic and attractive magnetic contributions to the CL force can be continuously tuned to form stable magnetic CL traps appearing as equilibria whose strength and range derive not from mechanical considerations but from spectral properties of quantum electromagnetic fluctuations.
+Here, we thus propose a novel experimentally feasible approach to manipulate nanorod pair complexes by suspending them in a magnetically diluted fluid, where the interplay between repulsive non-magnetic and attractive magnetic contributions to the CL force can be continuously tuned to form magnetic CL traps appearing as equilibria whose strength and range derive not from mechanical considerations but from spectral properties of quantum electromagnetic fluctuations.
 Figure 1: Color online: Two cylindrical nanorods, one composed of ZnO with permittivity $\varepsilon_{1}$ and permeability $\mu_{1}$, and the other of PS with permittivity $\varepsilon_{2}$ and permeability $\mu_{2}$, each having an identical radius $a$, are immersed in a magnetic liquid.
-The magnetic liquid is composed of toluene and it contains dispersed magnetite nanoparticles of diameter $D$ and volume fraction $\phi$ and it has effective permittivity $\varepsilon_{3}$ and permeability $\mu_{3}$.
+The magnetic liquid is composed of toluene and it contains dispersed gold coated (thickness $R_{\rm Au}$) magnetite nanoparticles of diameter $D$ and volume fraction $\phi$ and it has effective permittivity $\varepsilon_{3}$ and permeability $\mu_{3}$.
 The nanorods are separated by a center-to-center distance $R$.
@@
 We assume that the cylinders are infinitely long in the $z$ direction, and we will work in the quasi-static limit where the wavelength of light is much larger than any other length scale in the problem.
-The dispersion interaction in this system is obtained using scattering approach 25, where the interaction energy per unit length (for a cylinder with length $L$) in finite temperature $T$ is (an explicit derivation is given in the Supplementary Information),
+The dispersion interaction in this system is obtained using scattering approach NylandBrevik1994, where the interaction energy per unit length (for a cylinder with length $L$) in finite temperature $T$ is (an explicit derivation is given in Sec.
+(S.2) of the Supplementary Information),
 $G(a,R,T)=\frac{k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\,\ln\mathscr{D}(i\xi_{p})$ (1)
@@
 The above expression is given in Matsubara frequencies $\xi_{p}=2\pi k_{B}Tp/\hbar$, where $p$ is an integer.
-Here, $\mathscr{D}(i\xi_{p})$ is expressed as the determinant of the scattering matrices found in the “thin cylinder approximation” e.g., the separation distance between the cylinders is much larger than their radii ($R\gg a$).
-We find that (see Supplementary Information for details),
+Here, $\mathscr{D}(i\xi_{p})$ is expressed as the determinant of the scattering matrix of the interacting cylinders.
+In general, such a matrix is infinite, but significant simplifications can be made in the “thin cylinder” approximation where the separation distance between the cylinders is much larger than their radii ($R\gg a$).
+In this case, terms containing Bessel functionswatson1958 with $n\geq 2$ make negligible contributions (details in Sec.
+(S.2) in the Supplementary Information) as they rapidly decay with $R$.
+Thus, we find that (see Sec.
+(S.2) of the Supplementary Information),
 $\ln\mathscr{D}(i\xi_{p})=f_{0}K^{2}_{0}(\nu_{3}R)+f_{1}K^{2}_{1}(\nu_{3}R)+f_{2}\Big(K^{2}_{0}(\nu_{3}R)+K^{2}_{2}(\nu_{3}R)\Big)$ (2)
@@
 Here we are also interested if a dispersion interaction in excited state could be realized between two equal cylindrical nanoparticles.
-In order to gain an understanding into this issue, it is instructive to start by considering the corresponding analogous situation of two identical atoms; one in its excited state while the other in its ground state.
-In such a scenario, the interaction can be described as a superposition of a symmetric and an antisymmetric state 28.
-Among these, the only state that is of practical interest will be the antisymmetric one which is often very long-lived thus effectively holding the system as a trapped metastable state.
-As a result, a strong resonant dispersion interaction will be maintained between the atoms until (I) the atoms move sufficiently far apart or (II) a photon is emitted, which reduces most of the coherence between them.
 The interaction energy is due to the separation dependence of the excited energy, such that $E^{\text{res}}(R)=\hbar[\omega_{r}(R)-\omega_{r}(\infty)]$, where $\omega_{r}(R)$ represents the resonance frequency at interatomic separation $R$.
-The interaction energy of the excited state is 29, 30, $E^{\text{res}}(R)=s^{2}T(R|\omega_{j})\propto R^{-3},$ where $s$ is the magnitude of the transition dipole moment of the excitation and $T(R|\omega_{j})$ is the dipole–dipole coupling tensor (from the electromagnetic Green’s function) at frequency $\omega_{j}$ which accounts for how field from one oscillating dipole influences another at distance $R$ and it typically scales as $R^{-3}$ in the non-retarded limit.
-This can be compared with the ground state van der Waals interaction which decays with distance as $R^{-6}$31.
+The interaction energy of the excited state is McLachlan, Bostrom1, $E^{\text{res}}(R)=s^{2}T(R|\omega_{j})\propto R^{-3},$ where $s$ is the magnitude of the transition dipole moment of the excitation and $T(R|\omega_{j})$ is the dipole–dipole coupling tensor (from the electromagnetic Green’s function) at frequency $\omega_{j}$ which accounts for how field from one oscillating dipole influences another at distance $R$ and it typically scales as $R^{-3}$ in the non-retarded limit.
+This can be compared with the ground state van der Waals interaction which decays with distance as $R^{-6}$Parsegian2006.
 This analogy implies that two identical cylindrical nanoparticles can similarly form a long lasting antisymmetric excited state by virtue of a long-range resonant dispersion interaction.
-A corresponding expression for the retarded excited-state interaction energy, obtained semiclassically in the Supplementary Information, reads
+A corresponding expression for the retarded excited-state interaction energy, obtained semiclassically in Sec.
+(S.3) of the Supplementary Information, reads
 $G^{\text{res}}(a,R,T)$ $=-\frac{2k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\,(t_{0}+t_{1}+t_{2}+t_{3})\,K_{0}(\nu_{3}R)$ (6)
@@
 The entropic contribution ($p=0$) suggests that Eq.
-(S.
-63) provides $R^{-3}$ dependence that is much longer ranged than the ground state interaction between two equal thin cylindrical nanoparticles that is proportional to $R^{-5}$32.
+(6) provides $R^{-3}$ dependence that is much longer ranged than the ground state interaction between two equal thin cylindrical nanoparticles that is proportional to $R^{-5}$langbein1972van.
 It gives both magnetic and non-magnetic contributions.
 Notably, this previously neglected magnetic contribution in Eq.
-(S.
-63) has the same power law as its non-magnetic contribution.
+(6) has the same power law as its non-magnetic contribution.
 However, the significance of this contribution should be judged by the low temperature/short distance non-magnetic excited state interaction.
@@
 In the retarded limit additional effects potentially impact the interaction via real photon exchange.
-Ultimately, in the limit of large distances or high temperatures, the interaction is expected to be dominated by the non-magnetic and magnetic zero frequency contributions.
+Ultimately, in the limit of large distances or high temperatures, the interaction is expected to be dominated by the non-magnetic and magnetic zero frequency contributions [for more detail see Sec.
+(S.7) in the Supplementary Information].
 However, it is important to emphasize that when the interaction occurs in a fluid the excited state lifetime will be significantly reduced.
@@
 The nanoparticles we consider have large enough dimensions to approximate their effective dielectric functions with those of bulk materials.
-The dielectric functions at imaginary Matsubara frequencies are obtained through the Kronig–Kramer relation
+The dielectric functions at imaginary Matsubara frequencies are obtained through the Kramers–Kronig relation
 $\varepsilon(i\xi_{p})=1+\frac{2}{\pi}\int_{0}^{\infty}d\omega\,\frac{\omega\,\varepsilon^{\prime\prime}(\omega)}{\omega^{2}+\xi_{p}^{2}}.$ (8)
-We use available experimental data from the literature for polystyrene (PS) 33 nanoparticles.
-This reference is also used to provide the dielectric function for Toluene.
-For zinc oxide (ZnO) 34 nanorods we use results from density functional theory (DFT) that, together with additional materials information, are described in detail in the Supplementary Information.
+The geometries of the nanoparticles and nanorods are described within our theoretical model, whereas the material properties are macroscopic attributes.
+We use available room-temperature experimental data for PS as provided by van Zwol, et al..
+Zwol1.
+For ZnO we calculate the imaginary part of the dielectric function within the density functional theory (DFT) as implemented in the Vienna Ab initio Simulation Package.
+VASP1999 To correctly mimic the experimental band-gap energy of 3.4 eV, we employ the hybrid functional with 37.5% Fock exact exchange, which is a widely adopted parameter choice for this oxide.
+The $GW$-type pseudopotentials ensure better accuracy in describing high-frequency optical transitions.
+Convergence with respect to $k$-space mesh is considered.
+Crovetto2016 Further computational details, together with additional materials information, are described in Sec.
+(S.5) of the Supplementary Information.
 The optical properties of the magnetic fluid are estimated by taking into account the volume fraction of commercially available gold-coated magnetite particles in a toluene mixture.
-We use the mixing model presented by Chen et al.35 as,
+We use the mixing model presented by Chen et al.ChenEffPermiLayered1998 as,
 $\varepsilon_{3}=\frac{\varepsilon_{3,3}(1+2\Lambda_{3})}{1-\Lambda_{3}};\quad\Lambda_{3}=(\phi_{1}+\phi_{2})\frac{\varepsilon_{s}-\varepsilon_{3,3}}{\varepsilon_{s}+2\varepsilon_{3,3}},$ (9)
@@
 $\varepsilon_{s}=\frac{\varepsilon_{3,1}(1-2G)}{1+2G};\quad G=\frac{\varepsilon_{3,1}-\varepsilon_{3,2}}{2\varepsilon_{3,1}+\varepsilon_{3,2}}\frac{\phi_{2}}{\phi_{1}+\phi_{2}}.$ (10)
-Here, $\phi_{i}$ and $\varepsilon_{3,i}$ are the volume fraction and dielectric function of toluene ($i=3$), magnetite ($i=2$) and gold ($i=1$) with gold coating thickness $R_{\rm Au}$.
+Here, $\phi_{i}$ and $\varepsilon_{3,i}$ are the volume fraction and dielectric function of toluene ($i=3$), magnetite ($i=2$) and gold ($i=1$) for modeling the magnetic fluid where $D$ is the diameter of the magnetite core and the thickness $R_{\rm Au}$ of the gold coating, as illustrated in Fig.
+(1).
+The radius of the gold coated nanoparticles is defined as $R_{1}=D/2+R_{\rm Au}$.
 In the current work, we consider the case where the amount of magnetite is 5% and 7% ($\phi_{2}=0.05,0.07$).
-Hence, it is of interest to link $\phi_{1}=\phi_{2}\times[(R_{1}/D)^{3}-1]$ where $R_{1}=D+R_{\rm Au}$.
-We consider four typical cases where the average diameter ($D$) of the magnetite nanoparticles is 10 nm, 13 nm, 15 nm, and 17 nm.
-We use experimental data from the literature for toluene 33 at $T=298.15$ K.
-Gold coatings are taken to have bulk properties given in the literature 36.
+The corresponding volume fraction of the gold coating is therefore given by $\phi_{1}=\phi_{2}\times[(2R_{1}/D)^{3}-1]$.
+We consider four typical cases where the average diameter ($D$) of the magnetite nanoparticles is 10, 13, 15 and 17 nm.
+We use experimental data for toluene at $T=298.15$ K from the literature.
+Zwol1 The dielectric functions of magnetite and gold are from DFT calculations in our recent work, Carretero_PRB_2025_PhysRevB.111.085407 where room-temperature thermal electronic excitation was considered for the small-gap magnetite compound.
 In the main text, we report some important results that are obtained for pure magnetite particles (i.e.
-for $R_{1}=D$).
+for $R_{1}=D/2$).
 Most of the results obtained using gold-coating are included in Sec.
 (S.8) of the Supplementary Information.
-The size of the magnetite nanoparticles and gold-coated magnetite nanoparticle can be adjusted to control trapping behaviour for pairs of unequal nanorods.
+The size of the magnetite nanoparticles and gold-coated magnetite nanoparticle can be adjusted to control trapping behavior for pairs of unequal nanorods.
 […]
-Figure 2: (Color online): The parametrized dielectric functions on the imaginary axis as a function of the frequencies ${\xi}_{p}$ for PS, and magnetic toluene based on optical experiments given by van Zwol et al.33.
-For ZnO, we performed DFT calculation [see Sec.
-(S.5.1) in the Supplementary Information].
-The dielectric function for the magnetic fluid is calculated using Eq.
-(9) and data from Ref.
-36 with a volume fraction of $\phi=5\%$.
-For water (close to being a non-magnetic fluid) with a fraction $\phi$ of magnetite Klimchitskaya et al. estimated the permeability to be 17, 18
+Figure 2: (Color online): The dielectric functions on the imaginary axis as functions of frequency $\xi$.
+The parametrized data for PS and toluene are based on optical experiments given by van Zwol et al.Zwol1.
+The magnetic fluid consists of a mixture of toluene and magnetite particles, using Eq.
+(9) for a volume fraction of $\phi=5\%$.
+The response functions of gold and magnetite are DFT results presented in Ref.
+Carretero_PRB_2025_PhysRevB.111.085407, whereas the data for ZnO are from present DFT calculations.
+For water (close to being a non-magnetic fluid) with a fraction $\phi$ of magnetite Klimchitskaya et al. estimated the permeability to be klimchitskaya2019impact, VelichkoKlimchitskaya2020
 $\mu_{3}(0)=1+\frac{2\pi^{2}\phi}{9}\frac{M_{s}^{2}D^{3}}{k_{B}T}.$ (11)
-Toluene is only a little more magnetic ($\mu=0.99996\approx 1.00$) so within the level of accuracy in our work the permeability of both water and toluene-based mixtures can be estimated with the approximation used by Klimchitskaya and Mostepanenko for water-based mixtures 17, 18.
+Toluene is only a little more magnetic ($\mu=0.99996\approx 1.00$) so within the level of accuracy in our work the permeability of both water and toluene-based mixtures can be estimated with the approximation used by Klimchitskaya and Mostepanenko for water-based mixtures klimchitskaya2019impact, VelichkoKlimchitskaya2020.
 Using Eq.
@@
 Figure 3: (Color online:) The CL force per unit length between two cylindrical objects (ZnO and PS respectively) with the same radius $a=20$ nm as a function of center-to-center distance $R$ when immersed in a ferrofluid with 5% and 7% concentrations of magnetite nanoparticles of diameter 10, 13, 15 and 17 nm.
-We recently investigated the interaction between a planar PTFE surface and a spherical PS surface in the presence of a magnetic fluid 21.
+We recently investigated the interaction between a planar PTFE surface and a spherical PS surface in the presence of a magnetic fluid pal2025trapping.
 In this case both the non-retarded and retarded interaction could be tuned to be repulsive while zero frequency contributions to the CL interaction (both TM and TE) where attractive.
-This led us to conclude that it is possible to have a magnetic contribution to CL trapping 21.
+This led us to conclude that it is possible to have a magnetic contribution to CL trapping pal2025trapping.
 However, trapping still occurred when the magnetic effects were ignored (i.e.
 the magnetic permeability set to one).
-Figure 4: (Color online): The CL force per unit length between two ZnO and PS cylindrical objects was evaluated for volume fraction $\phi=5\%$ and a nanoparticle diameter $D=17$ nm.
+Figure 4: (Color online): The CL force per unit length between two ZnO and PS cylindrical objects was evaluated for volume fraction $\phi=5\%,7\%$ and a nanoparticle diameter $D=17$ nm.
 The terms “without gold” and “with gold” refer to magnetite nanoparticles without and with a gold coating, respectively.
@@
 Here the intermediate magnetic fluid has a dielectric function in between those of ZnO and PS.
-This indicates that the non-retarded and retarded CL interaction between a ZnO and PS nanorod will be repulsive 4.
+This indicates that the non-retarded and retarded CL interaction between a ZnO and PS nanorod will be repulsive Dzya.
 In Fig.
@@
 This dependence of trapping on the magnetic permeability has not been possible to achieve in the past.
-Notably, when the diameter of the magnetite nanoparticles decreases with fixed volume fraction of magnetite, or a weak magnetic field is applied, leading to reduced magnetic permeability of the fluid 37, the trapping occurs at increasingly larger nanorod pair separations.
-Ultimately, in the absence of magnetic permeability effects, there is no trapping at all and the nanorod complex is split into two separate nanorods.
+Notably, when the diameter of the magnetite nanoparticles decreases with fixed volume fraction of magnetite, or a weak magnetic field is applied, leading to reduced magnetic permeability of the fluid ZhangNature2024, the trapping occurs at increasingly larger nanorod pair separations.
+The trapping can also be controlled by keeping the diameter constant and varying the volume fraction.
+As illustrated in Fig.
+(5), the trend is that larger $\phi$ shifts the force minimum towards larger separations and making it more shallow.
+Ultimately, in the absence of magnetic permeability effects or very small volume fraction, there is no trapping at all and the nanorod complex is split into two separate nanorods.
+Figure 5: (Color online:) The CL force per unit length between two cylindrical objects (ZnO and PS respectively) with the same radius $a=20$ nm as a function of center-to-center distance $R$ when immersed in a ferrofluid with different concentrations of ($\phi$=0.01-0.15) magnetite nanoparticles of diameter 17 nm.
 Replacing the magnetite nanoparticles with commercially available gold-coated magnetite nanoparticles leads to an additional mechanism to tune the trapping distance as can be seen in Fig.
@@
 With changes in the gold coating the dielectric function of the background media changes and the trapping distance can be shifted.
-This framework suggests a new route for actively tuning the interacting forces between two unequal nanorods where a purely attractive potential can be driven into repulsive–attractive trapping potentials and even into regimes of complete repulsion.
+Non-magnetic Casimir forces have been exploited, both theoretically and experimentally, to produce Casimir trapping and non-adhesion, via repulsive forces, in NEMS and MEMSMunday2009, esteso2015nanolevitation, WoodsRevModPhys.88.045003, zhao2019stable, Woods2020, esteso2022effect, MundayReview2024, WilliamsonJPCL2025.
+We have focused in the current work on the novel results where only the zero-frequency TE term changes, as trapping occurs over a range of different $\phi$ values.
+The repulsion occurs when $\phi$ is tuned such that the dielectric function for the fluid is between the values of the dielectric functions of the interacting nanorods for a wide range of frequencies.
+Our work demonstrates that the trapping can be controlled by both changes in the size of the magnetic nanoparticles and the concentration in the fluid.
+The exact values for these quantities can be adjusted to achieve trapping for different separations and for different material combinations.
+The proposed framework suggests a novel route for actively tuning the interacting forces between two unequal nanorods where a purely attractive potential can be driven into repulsive–attractive trapping potentials and even into regimes of complete repulsion.
 By a careful choice of the material parameters, one can tailor finite nanosystems that give rise to reversible adhesive behavior, stable tunable trapping and on-demand destabilization of clusters.
@@
 Nevertheless, this restriction can be mitigated by applying the proximity force approximation (PFA), valid in the short-distance regime, where enhanced stability of the trapping potential is expected.
-Going forward, the framework we have developed here indicates some exciting prospects for future extension of tunable dispersion-force control to biomolecular materials.
+Going forward, the framework we have developed here indicates some exciting prospects for future extension of tunable dispersion-force controlMunday2009, esteso2015nanolevitation, WoodsRevModPhys.88.045003, zhao2019stable, Woods2020, esteso2022effect, MundayReview2024, WilliamsonJPCL2025 to biomolecular materials.
 In particular, these ideas could be extended to direct biopolymer complexation with artificial nanothreads via magnetic nanoparticles in biofluids with a suitable choice of materials.
-This has a great potential in a variety of future diagnostic and therapeutic nanomedicine research 38.
+This has a great potential in a variety of future diagnostic and therapeutic nanomedicine research CoeneMagneticNPJApplPhys2022.
 It also provides a route to field-controlled assembly in hybrid soft–magnetic nanostructures and could also find utility for biosensing, targeted delivery, reconfigurable nanoarchitectures, and beyond.
+Data availability: The code and data that support the findings of this study are available in Github.
 1 Acknowledgement
@@
 PLL/2023/4/016319.
-CP acknowledges the access to high-performance computing resources from NAISS, provided by NSC and PDC.
+CP acknowledges access to high-performance computing resources from NAISS, provided by NSC and PDC.
 LMW acknowledges financial support from the US Department of Energy under Grant No.
 DE-FG02-06ER46297.
-Supplementary Information
-Nanorod Pair Complexes Manipulated via Magnetic Casimir Forces
-S Pal , L. M. Woods , C. Persson , I. Brevik ,
-U. De Giovannini ,  and M. Boström
-*Corresponding authors: palsubhojit429@gmail.com, umberto.degiovannini@unipa.it, mathias.bostrom@ensemble3.eu
-S.1 Maxwell’s equations
-We start with Maxwell’s equations where there is no current density fluctuations ($J=0$) and charge density fluctuations $(\rho=0)$ written as 39
-$\nabla\cdot E=0\,\quad\quad\quad\nabla\cdot D=0\,(D=\varepsilon E)$ (S. 1)
-$\nabla\cdot B=0\,\,\,\quad\quad\quad\nabla\cdot B=0\,(B=\mu H)$
-$\nabla\times E=-\frac{\partial B}{\partial t}\quad\quad\quad\nabla\times E=-\frac{1}{c}\frac{\partial B}{\partial t}$
-$\nabla\times B=\mu_{0}\varepsilon_{0}\frac{\partial E}{\partial t}\quad\quad\quad\nabla\times H=\frac{1}{c}\frac{\partial D}{\partial t}$
-We can deduce a further equation from the last two expressions of the right side panel as,
-$\nabla\times(\nabla\times E)=-\frac{\mu}{c^{2}}\frac{\partial}{\partial t}(\nabla\times H)\implies\nabla^{2}E+\frac{\mu\varepsilon}{c^{2}}\frac{\partial^{2}E}{\partial t^{2}}=0$ (S. 2)
-Throughout the calculations, we consider the time dependency of the parameters $\sim e^{-i\omega t}$. After Fourier transform we obtain an expression of electric field $E$ and magnetic field $H$
-$\nabla^{2}E+\frac{\omega^{2}}{c^{2}}\mu\varepsilon E=0$ (S. 3)
-$\nabla^{2}H+\frac{\omega^{2}}{c^{2}}\mu\varepsilon H=0$ (S. 4)
-We proceed to consider an ansatz of Eq. (S. 3) and of (S. 4) as
-$E=E(r,\theta)\,e^{i(kz-\omega t)}\quad H=H(r,\theta)\,e^{i(kz-\omega t)}$ (S. 5)
-where $z$ axis is parallel ($||^{n}$) to the cylinders long axis. More compactly, we can write
-$\nabla^{2}_{r,\theta}E+u^{2}E=0\quad\&\quad\nabla^{2}_{r,\theta}H+u^{2}H=0$ (S. 6)
-where $\nabla^{2}_{r,\theta}$ is the two-dimensional Laplacian in polar coordinates and $u^{2}=\Big(\frac{\omega^{2}}{c^{2}}\mu(\omega)\varepsilon(\omega)-k^{2}\Big)$.
-A general solution of the magnetic field in Eq.
-(S.
-6) ($r<a$) can be obtained in cartesian coordinates for one cylinder placed at the origin as shown in Fig.
-(1) in the main text,
-$H^{\text{in}}_{z}=\sum_{n}B_{n}J_{n}(ur)e^{in\theta}$ (S. 7a)
-$H^{\text{in}}_{x}=\sum_{n}X_{n}J_{n}(ur)e^{in\theta}$ (S. 7b)
-$H^{\text{in}}_{y}=\sum_{n}Y_{n}J_{n}(ur)e^{in\theta}$ (S. 7c)
-Our main goal is to represent the radial and tangential components of magnetic field, $H^{\text{in}}_{r}$ and $H^{\text{in}}_{\theta}$ in terms of the cartesian components $H^{\text{in}}_{x}$ and $H^{\text{in}}_{y}$ in Eqs.
-(S.
-7b) and (S.
-7c).
-Following standard notation, we can easily write
-$H^{\text{in}}_{r}=H^{\text{in}}_{x}\cos(\theta)+H^{\text{in}}_{y}\sin(\theta)$ (S. 8)
-$H^{\text{in}}_{\theta}=-H^{\text{in}}_{x}\sin(\theta)+H^{\text{in}}_{y}\cos(\theta)$ (S. 9)
-Using Eqs. (S. 7b) and (S. 7c), we rewrite Eqs. (S. 8) and (S. 9) as,
-$H^{\text{in}}_{r}=\sum_{n}\Big(X_{n}\cos(\theta)+Y_{n}\sin(\theta)\Big)J_{n}(ur)e^{in\theta}$ (S. 10)
-$H^{\text{in}}_{\theta}=\sum_{n}\Big(Y_{n}\cos(\theta)-Y_{n}\sin(\theta)\Big)J_{n}(ur)e^{in\theta}$ (S. 11)
-With the help of $\cos(\theta)=\frac{e^{i\theta}+e^{-i\theta}}{2}$ and $\sin(\theta)=\frac{e^{i\theta}-e^{-i\theta}}{2i}$, the radial component of the magnetic field in Eq.
-(S.
-10) can be simplified as,
-$H^{\text{in}}_{r}=\sum_{n}\Bigg[\Big(\frac{X_{n}e^{i(n+1)\theta}-iY_{n}e^{i(n+1)\theta}}{2}\Big)J_{n}(ur)+\Big(\frac{X_{n}e^{i(n-1)\theta}+iY_{n}e^{i(n-1)\theta}}{2}\Big)J_{n}(ur)\Bigg]$ (S. 12)
-Replace $n$ by n1n-1 in the first term and $n$ by $n+1$ in the second term of Eq. (S. 12), yields
-$H^{\text{in}}_{r}=\sum_{n}\Big[W_{n}J_{n-1}(ur)+Z_{n}J_{n+1}(ur)\Big]e^{in\theta}$ (S. 13)
-where $W_{n}=\frac{1}{2}\Big(X_{n-1}-iY_{n-1}\Big)$ and $Z_{n}=\frac{1}{2}\Big(X_{n+1}+iY_{n+1}\Big)$.
-Similarly we obtain a simplified expression for tangential component $H^{\text{in}}_{\theta}$ in Eq.
-(S.
-11) as,
-$H^{\text{in}}_{\theta}=i\sum_{n}\Big[W_{n}J_{n-1}(ur)-Z_{n}J_{n+1}(ur)\Big]e^{in\theta}$ (S. 14)
-As $B_{n}$, $W_{n}$ and $Z_{n}$ are not independent parameters, we further connect them using the Maxwell equation $\nabla\cdot H=0$, which in cylindrical coordinate can be expressed as,
-$\frac{1}{r}\frac{\partial}{\partial r}(rH_{r})+\frac{1}{r}\frac{\partial H_{\theta}}{\partial\theta}+ikH_{z}=0$ (S. 15)
-After substituting the expressions for $H_{r}$, $H_{\theta}$ and $H_{z}$ into Eq. (S. 15), we obtain
-$\frac{1}{ur}\Big[W_{n}J_{n-1}+Z_{n}J_{n+1}-nW_{n}J_{n-1}+nW_{n}J_{n+1}\Big]+W_{n}J^{\prime}_{n-1}+Z_{n}J^{\prime}_{n+1}+\frac{ik}{u}B_{n}J_{n}=0$ (S. 16)
-Now we can use the recurrence relation of ordinary Bessel functions $J_{n}$, $J^{\prime}_{n+1}(x)=\frac{1}{2}[J_{n-1}(x)-J_{n+1}(x)]$ , $\forall x$, that yields a compact relation between the coefficients $B_{n}$, $W_{n}$ and $Z_{n}$ as,
-$W_{n}+Z_{n}+\frac{ik}{u}B_{n}=0$ (S. 17)
-We know the $z$ component of electric field can be presented as,
-$E_{z}=\sum_{n}A_{n}J_{n}(ur)e^{in\theta}$ (S. 18)
-which is connected to the radial and tangential components of magnetic field $H_{r}$ and $H_{\theta}$ by another Maxwell equation
-$(\nabla\times H)_{z}=-\frac{i\omega\varepsilon}{c}E_{z}$ (S. 19)
-Again in cylindrical coordinate, Eq. (S. 19) can be written as
-$\frac{1}{r}\Big(H_{\theta}-\frac{\partial H_{r}}{\partial\theta}\Big)+\frac{\partial H_{\theta}}{\partial r}=-\frac{i\omega\varepsilon}{c}E_{z}$ (S. 20)
-Similarly by substituting the expressions of $H_{r}$, $H_{\theta}$ and $E_{z}$ into Eq.
-(S.
-20), we extract a relation between the coefficients $A_{n}$, $W_{n}$ and $Z_{n}$ after using the recurrence relation of ordinary Bessel functions,
-$W_{n}+Z_{n}=\frac{\omega\varepsilon}{cu}A_{n}$ (S. 21)
-Eqs.
-(S.
-17) and (S.
-21) determine the coefficients of $H_{r}$ and $H_{\theta}$ in terms of $A_{n}$ and $B_{n}$, the coefficients of the $z$ component of electric and magnetic field.
-After solving these two equations, we obtain
-$W_{n}=\frac{1}{2}\Big(\frac{\omega\varepsilon}{cu}A_{n}+\frac{ik}{u}B_{n}\Big)\quad\&\quad Z_{n}=\frac{1}{2}\Big(\frac{\omega\varepsilon}{cu}A_{n}-\frac{ik}{u}B_{n}\Big)$ (S. 22)
-These expressions of $W_{n}$ and $Z_{n}$ in Eq. (S. 22) lead to a set of normalized solutions for some of the components of $E$ and $H$ inside the cylindrical region from Eq. (S. 6) as,
-$E^{\text{in}}_{z}=\sum_{n=-\infty}^{\infty}A_{n}\frac{J_{n}(u_{1}r)}{J_{n}(u_{1}a)}e^{in\theta}$ (S. 23a)
-$H^{\text{in}}_{z}=\sum_{n=-\infty}^{\infty}B_{n}\frac{J_{n}(u_{1}r)}{J_{n}(u_{1}a)}e^{in\theta}$ (S. 23b)
-$H^{\text{in}}_{r}=\sum_{n=-\infty}^{\infty}\frac{1}{u_{1}}\Bigg(A_{n}\frac{n\omega\varepsilon_{1}}{cu_{1}r}\frac{J_{n}(u_{1}r)}{J_{n}(u_{1}a)}+B_{n}ik\frac{J^{\prime}_{n}(u_{1}r)}{J_{n}(u_{1}a)}\Bigg)e^{in\theta}$ (S.
-23c)
-$H^{\text{in}}_{\theta}=\sum_{n=-\infty}^{\infty}\frac{i}{u_{1}}\Bigg(A_{n}\frac{\omega\varepsilon_{1}}{c}\frac{J^{\prime}_{n}(u_{1}r)}{J_{n}(u_{1}a)}+B_{n}\frac{ikn}{u_{1}r}\frac{J_{n}(u_{1}r)}{J_{n}(u_{1}a)}\Bigg)e^{in\theta},$ (S.
-23d)
-where $u^{2}_{1}(\omega)=\Big(\frac{\omega^{2}}{c^{2}}\mu_{1}(\omega)\varepsilon_{1}(\omega)-k^{2}\Big)$, ‘1’ denotes the coordinates of cylinder 1 and $J_{n}$ is the ordinary Bessel function of first kind of order $n$ in standard mathematical notation.
-Similarly, the corresponding solutions outside of cylinder 1 are,
-$E^{\text{ext}}_{z}=\sum_{n=-\infty}^{\infty}\Bigg(C_{n}\frac{H_{n}(u_{3}r)}{H_{n}(u_{3}a)}+F_{n}\frac{J_{n}(u_{3}r)}{J_{n}(u_{3}a)}\Bigg)e^{in\theta}$ (S. 24a)
-$H^{\text{ext}}_{z}=\sum_{n=-\infty}^{\infty}\Bigg(D_{n}\frac{H_{n}(u_{3}r)}{H_{n}(u_{3}a)}+G_{n}\frac{J_{n}(u_{3}r)}{J_{n}(u_{3}a)}\Bigg)e^{in\theta}$ (S. 24b)
-$H^{\text{ext}}_{r}=\sum_{n=-\infty}^{\infty}\frac{1}{u_{3}}\Bigg(C_{n}\frac{n\omega\varepsilon_{3}}{cu_{3}r}\frac{H_{n}(u_{3}r)}{H_{n}(u_{3}a)}+D_{n}ik\frac{H^{\prime}_{n}(u_{3}r)}{H_{n}(u_{3}a)}+F_{n}\frac{n\omega\varepsilon_{3}}{cu_{3}r}\frac{J_{n}(u_{3}r)}{J_{n}(u_{3}a)}+G_{n}ik\frac{J^{\prime}_{n}(u_{3}r)}{J_{n}(u_{3}a)}\Bigg)e^{in\theta}$ (S.
-24c)
-$H^{\text{ext}}_{\theta}=\sum_{n=-\infty}^{\infty}\frac{i}{u_{3}}\Bigg(C_{n}\frac{\omega\varepsilon_{3}}{c}\frac{H^{\prime}_{n}(u_{3}r)}{H_{n}(u_{3}a)}+D_{n}\frac{ikn}{u_{3}r}\frac{H_{n}(u_{3}r)}{H_{n}(u_{3}a)}+F_{n}\frac{\omega\varepsilon_{3}}{c}\frac{J^{\prime}_{n}(u_{3}r)}{J_{n}(u_{3}a)}+G_{n}\frac{ikn}{u_{3}r}\frac{J_{n}(u_{3}r)}{J_{n}(u_{3}a)}\Bigg)e^{in\theta}$ (S.
-24d)
-where $u^{2}_{3}(\omega)=\Big(\frac{\omega^{2}}{c^{2}}\mu_{3}(\omega)\varepsilon_{3}(\omega)-k^{2}\Big)$ and $H_{n}$ is the Hankel function of first kind of order $n$ in standard mathematical notation.
-Here ‘3’ denotes the background media where the second terms of Eqs.
-(S.
-24a) and (S.
-24b) are required to satisfy the boundary conditions at the interfaces due to the presence of cylinder 2.
-All coefficients $A_{n}$, $B_{n}$, $C_{n}$, $D_{n}$, $F_{n}$ and $G_{n}$ are connected via boundary conditions at the edge of the cylinders $r=a$.
-It may be useful to explore in some more detail the physical content of these expansions.
-The reason why only the ordinary Bessel functions $J_{n}$ occur inside the cylindrical region, is that the fields are required to be finite at the center $r=0$.
-In the outside region, the coefficients $C_{n}$ and $D_{n}$ describe the outward moving wave, while the $F_{n}$ and $G_{n}$ coefficients $F_{n}$ capture the remaining fields.
-Here, $F_{n}$ and $G_{n}$ correspond to the external response to the incident wave.
-In all cases, the Bessel functions are solutions to the governing field equation, while the occurrence of two integration constants reflects the second order property of this equation.
-The coefficients $F_{n}$ and $G_{n}$ represent the incident wave, whereas the coefficients $C_{n}$ and $D_{n}$ describe the reflected wave, which are the external responses to the incident wave.
-To determine these coefficients, it is necessary to ensure the continuity of $E_{z}$, $H_{z}$, $H_{\theta}$, and $\mu H_{r}$.
-These boundary conditions yield the following expressions,
-$A_{n}=C_{n}+F_{n}\quad\&\quad B_{n}=D_{n}+G_{n}$ (S. 25)
-$-\frac{nk}{u^{2}_{1}a}B_{n}+\frac{i\omega\varepsilon_{1}}{cu_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{1}a)}A_{n}=\frac{i\omega\varepsilon_{3}}{cu_{3}}\frac{H^{\prime}_{n}(u_{3}a)}{J_{n}(u_{3}a)}C_{n}+\frac{i\omega\varepsilon_{3}}{cu_{3}}\frac{J^{\prime}_{n}(u_{3}a)}{J_{n}(u_{3}a)}F_{n}-\frac{nk}{u^{2}_{3}a}D_{n}-\frac{nk}{u^{2}_{3}a}G_{n}$ (S.
-26)
-$\mu_{1}\Bigg[\frac{n\omega\varepsilon_{1}}{cu^{2}_{1}a}A_{n}+\frac{ik}{u_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{1}a)}B_{n}\Bigg]=\mu_{3}\Bigg[\frac{n\omega\varepsilon_{3}}{cu^{2}_{3}a}C_{n}+\frac{ik}{u_{3}}\frac{H^{\prime}_{n}(u_{3}a)}{H_{n}(u_{3}a)}D_{n}+\frac{n\omega\varepsilon_{3}}{cu^{2}_{3}a}F_{n}+$ (S.
-27)
-$\frac{ik}{u_{3}}\frac{J^{\prime}_{n}(u_{3}a)}{J_{n}(u_{3}a)}G_{n}\Bigg]$
-where primed Bessel functions are derivative with respect to the radial coordinates.
-Now we can easily eliminate the $A_{n}$ and $B_{n}$ from Eqs.
-(S.
-26) and (S.
-27) using Eq.
-(S.
-25), then we express $C_{n}$ and $D_{n}$ in terms of $F_{n}$ and $G_{n}$ in a matrix form as,
-$\begin{pmatrix}C_{n}\\ D_{n}\end{pmatrix}=\begin{pmatrix}Q^{11}_{n}&Q^{12}_{n}\\ Q^{21}_{n}&Q^{22}_{n}\end{pmatrix}\begin{pmatrix}F_{n}\\ G_{n}\end{pmatrix}$ (S. 28)
-where the matrix elements are
-$Q^{11}_{n}=\Bigg(\frac{\alpha\delta^{\prime}-\alpha^{\prime}\delta}{\alpha\beta^{\prime}-\alpha^{\prime}\beta}\Bigg)\quad\&\quad Q^{12}_{n}=\Bigg(\frac{\alpha^{\prime}\gamma-\alpha\gamma^{\prime}}{\alpha\beta^{\prime}-\alpha^{\prime}\beta}\Bigg),$ (S.
-29)
-$Q^{21}_{n}=\Bigg(\frac{\beta\delta^{\prime}-\beta^{\prime}\delta}{\alpha\beta^{\prime}-\alpha^{\prime}\beta}\Bigg)\quad\&\quad Q^{22}_{n}=\Bigg(\frac{\beta^{\prime}\gamma-\beta\gamma^{\prime}}{\alpha\beta^{\prime}-\alpha^{\prime}\beta}\Bigg)$
-and
-$\gamma=-\alpha=\frac{kn}{a}\Big(\frac{1}{u^{2}_{1}}-\frac{1}{u^{2}_{3}}\Big),\quad\quad\beta^{\prime}=-\delta^{\prime}=\frac{n\omega}{ca}\Big(\frac{\varepsilon_{3}\mu_{3}}{u^{2}_{3}}-\frac{\varepsilon_{1}\mu_{1}}{u^{2}_{1}}\Big)$ (S.
-30)
-$\alpha^{\prime}=ik\Big(\frac{\mu_{1}}{u_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{1}a)}-\frac{\mu_{3}}{u_{3}}\frac{H^{\prime}_{n}(u_{3}a)}{H_{n}(u_{3}a)}\Big),\quad\gamma^{\prime}=ik\Big(\frac{\mu_{3}}{u_{3}}\frac{J^{\prime}_{n}(u_{3}a)}{J_{n}(u_{3}a)}-\frac{\mu_{1}}{u_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{1}a)}\Big)$
-$\beta=\frac{i\omega}{c}\Big(\frac{\varepsilon_{3}}{u_{3}}\frac{H^{\prime}_{n}(u_{3}a)}{H_{n}(u_{3}a)}-\frac{\varepsilon_{1}}{u_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{1}a)}\Big),\quad\delta=\frac{i\omega}{c}\Big(\frac{\varepsilon_{1}}{u_{1}}\frac{J^{\prime}_{n}(u_{1}a)}{J_{n}(u_{3}a)}-\frac{\varepsilon_{3}}{u_{3}}\frac{J^{\prime}_{n}(u_{3}a)}{J_{n}(u_{3}a)}\Big)$
-The above expressions completely determine the response of a single cylinder with radius $a$ to an arbitrary incident electromagnetic wave.
-For nonmagnetic medium, the matrix in Eq.
-(S.
-28) agrees with previously found results by Mitchell et al.23.
-At this point we consider the case when both the cylinders are present in the background media.
-It is noted that the arbitrary incident waves with coefficients $F_{n}$ and $G_{n}$ in the first cylinder are linked to a reflected wave coming from the second cylinder.
-The $z$ components of the reflected fields from the second cylinder can be written as,
-$E_{z}=\sum_{n=-\infty}^{\infty}C^{\prime}_{n}\frac{H_{n}(u_{3}r_{2})}{H_{n}(u_{3}a)}e^{in\theta_{2}}$ (S. 31)
-$H_{z}=\sum_{n=-\infty}^{\infty}D^{\prime}_{n}\frac{H_{n}(u_{3}r_{2})}{H_{n}(u_{3}a)}e^{in\theta_{2}}$
-To correlate between the incident waves onto the first cylinder and the reflected waves from the second cylinder we need to express $F_{n}$ and $G_{n}$ in terms of $C^{\prime}_{n}$ and $D^{\prime}_{n}$.
-The general strategy to do this is as follows: first we need to express the reflected fields of the second cylinder in the coordinates of the first cylinder ($r,\theta,z$) using Graf’s addition theorem 40 for Hankel function,
-$H_{n}(u_{3}r_{2})\,e^{in\theta_{2}}=\sum_{m=-\infty}^{\infty}H_{m+n}(u_{3}R)\,J_{m}(u_{3}r)\,e^{im\theta}$ (S. 32)
-Hence the magnetic field in Eq. (15) can be written as
-$H_{z}=\sum_{m=-\infty}^{\infty}\frac{J_{m}(u_{3}r)}{J_{m}(u_{3}a)}J_{m}(u_{3}a)\sum_{n=-\infty}^{\infty}D^{\prime}_{n}\frac{H_{m+n}(u_{3}R)}{H_{n}(u_{3}a)}e^{im\theta}$ (S. 33)
-We can now easily identify the coefficients of $e^{im\theta}$ with the coefficient $G_{m}$ by comparing Eqs. (S. 24b) and (S. 33) that yields
-$G_{m}=J_{m}(u_{3}a)\sum_{n=-\infty}^{\infty}D^{\prime}_{n}\frac{H_{m+n}(u_{3}R)}{H_{n}(u_{3}a)}$ (S. 34)
-Similarly we obtain a further equation for the coefficient $F_{m}$ of the $z$ component of electric field as,
-$F_{m}=J_{m}(u_{3}a)\sum_{n=-\infty}^{\infty}C^{\prime}_{n}\frac{H_{m+n}(u_{3}R)}{H_{n}(u_{3}a)}$ (S. 35)
-Using these expressions, Eq. (S. 28) is written as,
-$\begin{pmatrix}C_{m}\\ D_{m}\end{pmatrix}=\underbrace{\sum_{n}Q_{mn}\begin{pmatrix}Q^{11}_{m}&Q^{12}_{m}\\ Q^{21}_{m}&Q^{22}_{m}\end{pmatrix}}_{M_{mn}}\begin{pmatrix}C^{\prime}_{n}\\ D^{\prime}_{n}\end{pmatrix}$ (S.
-36)
-where
-$Q_{mn}=J_{m}(u_{3}a)\frac{H_{m+n}(u_{3}R)}{H_{n}(u_{3}a)}$ (S. 37)
-Now let’s consider $\widetilde{X}$ as an infinite column matrix where the matrix elements are   $\cdots C_{m},\,D_{m}\cdots$ and a infinite matrix $\widetilde{M}$ with corresponding matrix elements are $M_{mn}$ and another infinite column matrix $\widetilde{X}^{\prime}$ whose elements are   $\cdots C^{\prime}_{n},\,D^{\prime}_{n}\cdots$.
-A compact representation of Eq.
-(S.
-36) is,
-$\widetilde{X}=\widetilde{M}\widetilde{X}^{\prime}$ (S. 38)
-A similar type of expression exists when the interaction is viewed from the second cylinder,
-$\widetilde{X}^{\prime}=\widetilde{N}\widetilde{X}$ (S. 39)
-where $\widetilde{N}$ can be deduced from $\widetilde{M}$ upon replacing $u_{1}$ by $u_{2}$, $u^{2}_{2}(\omega)=\Big(\frac{\omega^{2}}{c^{2}}\mu_{2}(\omega)\varepsilon_{2}(\omega)-k^{2}\Big)$.
-From combing the above equations we are able to obtain the dispersion relation of the surface modes:
-$\mathscr{D}(\omega)\equiv\mathrm{Det}(\mathbb{I}-\widetilde{\Omega})=0;\quad\widetilde{\Omega}=\widetilde{M}\widetilde{N}$ (S. 40)
-Hence the free energy of interaction per unit length (for a cylinder with length $L$) in terms of the dispersion relation is given by
-$G(a,R,T)=\frac{k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\,\ln\mathscr{D}(i\xi_{p})$ (S. 41)
-where the prime on the summation means the term for which $p=0$ must be weighted by $1/2$ and the Matsubara frequency $\xi_{p}=2\pi kTp/\hbar$, $p$ is an integer.
-S.2 Dispersion interaction for thin cylinders
-Though the expression for the dispersion relation in Eq.
-(S.
-41) defines the full solution of our problem, it is very cumbersome and of no practical advantage in its original form.
-For more useful and transparent solutions we consider the special case of “thin cylinder approximation” in the limit of $a\ll R$.
-There exists here a physical lower limit on the magnitude of $a$: it must be so large that classical fluid mechanics is applicable.
-Usually, a lower limit of about 100 nm should be more than sufficient.
-One must observe that in the limit of very thin cylindrical structures, e.g.
-DNA strings, one must use a quantum mechanical formalism.
-To reduce the complexity of our solutions, we followed a method developed for finding the non-retarded interaction between cylinders 23.
-The main contribution comes from the range of axial wave vectors, $k\lesssim 1/R$, which along with the thin cylinder approximation $a/R\ll 1$ leads to a further approximation $|u_{\alpha}a|\ll 1,(\alpha=1,2,3)$ ($|\xi a/c|\ll 1$).
-In this region, the Bessel functions are expanded to a leading order term for its small arguments $u_{\alpha}a$ and $u_{\alpha}R$ and the approximated matrix elements $M_{mn}$ can be written as,
-$M_{mn}\simeq\Big(\frac{a}{R}\Big)^{m+n}\frac{(m+n-1)!}{m!(n-1)!}<\Big(\frac{a}{R}\Big)^{m+n}2^{m+n}$ (S. 42)
-This shows that with increasing $m$ and $n$, these matrix elements decrease rapidly. On the other hand when $k>1/R$, using Matsubara frequencies, we find
-$u^{2}_{\alpha}(i\xi_{p})=-\Big(\frac{\xi^{2}_{p}}{c^{2}}\mu_{\alpha}(i\xi_{p})\varepsilon_{\alpha}(i\xi_{p})+k^{2}\Big)$ (S. 43)
-At this point, it is convenient to introduce a new variable $\nu_{\alpha}=e^{i\pi/2}u_{\alpha}$ leading to a new relation for the modified Bessel function,
-$H_{n}(xe^{i\pi/2})=-\frac{2i}{\pi}\,e^{-in\pi/2}K_{n}(x),\quad-\pi<\mathrm{arg}(x)<\pi/2$ (S. 44)
-Hence the matrix elements $M_{mn}$ can be connected with the modified Bessel function, using the new variable $\nu_{\alpha}$,
-$M_{mn}\propto H_{m+n}(u_{3}R)\propto K_{m+n}(\nu_{3}R)$ (S. 45)
-Therefore, in the region $kR\gg 1$ i.e., $|\nu_{3}R|\gg 1$, the matrix elements become
-$M_{mn}\simeq\sqrt{\frac{1}{\nu_{3}R}}\,e^{-\nu_{3}R}$ (S. 46)
-The above asymptotic expression, shows that in this limit the matrix elements are finite and they are independent of $m$ and $n$.
-Thus, we obtain finite contributions in the long and short separation limits.
-From Eqs.
-(S.
-42) and (S.
-46), it is very clear that the dominant effect can only be found in lower values of $m$ and $n$; i.e., $m,n=0,\pm 1$.
-The leading term which we require is then obtained by expanding the logarithmic term in Eq.
-(S.
-41) as,
-[…]
-[…]
-$\simeq-[M_{00}\,N_{00}+2M_{11}\,N_{11}+2M_{10}\,N_{01}+2M_{01}\,N_{10}+2M_{1-1}\,N_{-11}]$
-Using Eqs. (S. 36) and (S. 39), the matrix elements can be written down as,
-$-[M_{00}\,N_{00}+2M_{11}\,N_{11}+2M_{10}\,N_{01}+2M_{01}\,N_{10}+2M_{1-1}\,N_{-11}]=-Q_{00}S_{00}(Q^{11}_{0}S^{11}_{0}+Q^{22}_{0}S^{22}_{0})$ (S. 48)
-$-2Q_{01}S_{10}(Q^{11}_{0}S^{11}_{1}+Q^{22}_{0}S^{22}_{1})-2Q_{10}S_{01}(Q^{11}_{1}S^{11}_{0}+Q^{22}_{1}S^{22}_{0})-2(Q_{11}S_{11}+Q_{1,-1}S_{-1,1})$
-$(Q^{11}_{1}S^{11}_{1}+Q^{12}_{1}S^{21}_{1}+Q^{21}_{1}S^{12}_{1}+Q^{22}_{1}S^{22}_{1})$
-where the functions $S$ are determined from the functions $Q$ by replacing $\varepsilon_{1}$ and $\mu_{1}$ by $\varepsilon_{2}$ and $\mu_{2}$ respectively.
-It is worth mentioning that when the order of the ordinary Bessel function and Hankel function is not equal zero, then they can be approximated as,
-$\frac{J^{\prime}_{n}(x)}{J_{n}(x)}\simeq\frac{n}{x}\quad\quad\frac{H^{\prime}_{n}(x)}{H_{n}(x)}\simeq-\frac{n}{x}$ (S. 49)
-Using limiting forms of Bessel functions and Hankel functions, we list the values of matrix elements in Eq. (S. 36) in the following way:
-$Q^{11}_{n}$ $=-\Bigg[\frac{\varepsilon_{1}(\mu_{1}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{1})}{u^{2}_{1}}-\frac{2\varepsilon_{3}\mu_{3}}{u^{2}_{3}}\Bigg]\frac{u^{2}_{1}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})},\quad n\neq 0$ (S.
-50)
-$=\Bigg[\frac{\varepsilon_{3}-\varepsilon_{1}}{2\varepsilon_{3}}\Bigg]a^{2}u^{2}_{3}\ln(u_{3}a),\quad n=0$
-$Q^{12}_{n}$ $=\Bigg[\frac{ikc\mu_{3}u^{2}_{1}}{\omega(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})}\Bigg]\Bigg[\frac{1}{u^{2}_{1}}-\frac{1}{u^{2}_{3}}\Bigg],\quad n\neq 0$ (S. 51)
-$=0,\quad n=0$
-$Q^{21}_{n}$ $=\frac{i\omega}{kc}\Bigg[\frac{\varepsilon_{1}\mu_{1}}{u^{2}_{1}}-\frac{\varepsilon_{3}\mu_{3}}{u^{2}_{3}}\Bigg]\frac{2\varepsilon_{3}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})}\,u^{2}_{1},\quad n\neq 0$ (S.
-52)
-$=0,\quad n=0$
-$Q^{22}_{n}$ $=-\Bigg[\frac{\varepsilon_{3}(\mu_{1}+\mu_{3})+\varepsilon_{1}(\mu_{1}-\mu_{3})}{u^{2}_{1}}-\frac{2\varepsilon_{3}\mu_{3}}{u^{2}_{3}}\Bigg]\frac{u^{2}_{1}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})},\quad n\neq 0$ (S.
-53)
-$=\Bigg[\frac{\mu_{3}-\mu_{1}}{2\mu_{3}}\Bigg]a^{2}u^{2}_{3}\ln(u_{3}a),\quad n=0$
-when $x\to 0$, $K_{0}(x)\simeq-\ln(x)$. Therefore the elements corresponding to Eq. (S. 37) are expressed as,
-$\begin{split}Q_{00}=-\frac{i\pi}{2}\frac{H_{0}(u_{3}R)}{\ln(u_{3}a)},\quad\&\quad Q_{01}=ia\pi\Big(\frac{u_{3}}{2}\Big)H_{0}(u_{3}R)\\ Q_{10}=-a\Big(\frac{i\pi u_{3}}{4}\Big)\frac{H_{0}(u_{3}R)}{\ln(u_{3}a)},\quad\&\quad Q_{11}=a^{2}\Big(\frac{i\pi u^{2}_{3}}{4}\Big)H_{2}(u_{3}R)\\ Q_{1,-1}=-ia^{2}\pi\Big(\frac{u_{3}}{2}\Big)^{2}H_{0}(u_{3}R)\end{split}$ (S.
-54)
-and
-$Q_{11}=-Q_{-1,-1};\quad Q_{-1,1}=Q_{1,-1}$ (S. 55)
-$Q_{01}=Q_{0,-1};\quad\quad Q_{10}=Q_{-1,0}$
-Furthermore, using Eqs. (S. 48) and (S. 50)-(S. 55), Eq. (S. 47) can be written as,
-$\ln\mathscr{D}(i\xi_{p})=f_{0}K^{2}_{0}(\nu_{3}R)+f_{1}K^{2}_{1}(\nu_{3}R)+f_{2}\Big(K^{2}_{0}(\nu_{3}R)+K^{2}_{2}(\nu_{3}R)\Big)$ (S. 56)
-where $K_{0},K_{1}$ and $K_{2}$ are modified Bessel functions of zeroth, first and second orders respectively, and the coefficients are as follows:
-$f_{0}=-\frac{a^{4}}{4}\Bigg[\frac{(\varepsilon_{3}-\varepsilon_{1})(\varepsilon_{3}-\varepsilon_{2})}{\varepsilon^{2}_{3}}+\frac{(\mu_{3}-\mu_{1})(\mu_{3}-\mu_{2})}{\mu^{2}_{3}}\Bigg]\nu^{4}_{3}$ (S.
-57)
-$f_{1}=-a^{4}\Bigg\{\frac{\nu^{2}_{1}\nu^{4}_{3}}{(\varepsilon_{3}+\varepsilon_{1})(\mu_{3}+\mu_{1})}\Bigg[\frac{(\varepsilon_{3}-\varepsilon_{2})}{\varepsilon_{3}}\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{1}(\mu_{1}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{1})}{2\nu^{2}_{1}}\Bigg)+\frac{(\mu_{3}-\mu_{2})}{\mu_{3}}$ (S.
-58)
-$\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{1}+\mu_{3})+\varepsilon_{1}(\mu_{1}-\mu_{3})}{2\nu^{2}_{1}}\Bigg)\Bigg]+\frac{\nu^{2}_{2}\nu^{4}_{3}}{(\varepsilon_{3}+\varepsilon_{2})(\mu_{3}+\mu_{2})}\Bigg[\frac{(\varepsilon_{3}-\varepsilon_{1})}{\varepsilon_{3}}$
-$\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{2}(\mu_{2}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{2})}{2\nu^{2}_{2}}\Bigg)+\frac{(\mu_{3}-\mu_{1})}{\mu_{3}}\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{2}+\mu_{3})+\varepsilon_{2}(\mu_{2}-\mu_{3})}{2\nu^{2}_{2}}\Bigg)\Bigg]\Bigg\}$
-$f_{2}=-\frac{a^{4}\nu^{2}_{1}\nu^{2}_{2}\nu^{4}_{3}}{(\varepsilon_{3}+\varepsilon_{1})(\varepsilon_{3}+\varepsilon_{2})(\mu_{3}+\mu_{1})(\mu_{3}+\mu_{2})}\Bigg[\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{1}(\mu_{1}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{1})}{2\nu^{2}_{1}}\Bigg)$ (S.
-59)
-$\times\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{2}(\mu_{2}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{2})}{2\nu^{2}_{2}}\Bigg)+\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{1}}\Bigg)\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{2}\mu_{2}}{\nu^{2}_{2}}\Bigg)$
-$+\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{2}}\Bigg)\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{1}\mu_{1}}{\nu^{2}_{1}}\Bigg)+\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{1}+\mu_{3})+\varepsilon_{1}(\mu_{1}-\mu_{3})}{2\nu^{2}_{1}}\Bigg)\times$
-$\Bigg(\frac{\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}-\frac{\varepsilon_{3}(\mu_{2}+\mu_{3})+\varepsilon_{2}(\mu_{2}-\mu_{3})}{2\nu^{2}_{2}}\Bigg)\Bigg]$
-S.3 Excited state interaction derivation
-In the case of two identical thin cylindrical nanoparticles, if a long-lived antisymmetric state exists it results in a long-ranged excited state interaction.
-In general, this would formally be given by 26,
-$G^{\text{res}}(a,R,T)=\frac{2k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\ln[{\mathrm{Det}(\mathbb{I}-M)}]=-\frac{2k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\,\mathrm{Tr}[M_{mn}]$ (S.
-60)
-We are only interested in the first few values of $m,n=0,\pm 1$ in $M_{mn}$, as matrix elements decrease rapidly with increasing $m$ and $n$ (the follow an exponential decay in thin cylinder approximation and $kR\gg 1$).
-The trace of the matrix elements $M_{mn}$ can be written from its definition as,
-$\mathrm{Tr}[M_{mn}]=\sum_{m,n=-1}^{1}Q_{mn}\big(Q^{11}_{m}+Q^{22}_{m}\big)$ (S. 61)
-Ignoring the imaginary terms, we can write down all the matrix elements explicitly,
-$\mathrm{Tr}[M_{mn}]=2Q^{11}_{-1}Q_{1,-1}+2Q^{22}_{-1}Q_{1,-1}+Q_{00}\,Q^{11}_{0}+Q_{00}\,Q^{22}_{0}$ (S. 62)
-A relevant expression for the fully retarded excited-state interaction using Eqs.
-(S.
-50)-(S.
-55) and $H_{0}(u_{3}R)=-\frac{2i}{\pi}K_{0}(\nu_{3}R),\,\,-\pi<\mathrm{arg}(u_{3})<\pi/2$, $\nu^{2}_{\alpha}=-u^{2}_{\alpha},\quad\alpha=1,2,3$ is
-$G^{\text{res}}(a,R,T)$ $=-\frac{2k_{B}T}{\pi}\sum_{p=0}^{\infty}{}^{\prime}\int_{0}^{\infty}dk\,(t_{0}+t_{1}+t_{2}+t_{3})\,K_{0}(\nu_{3}R)$ (S. 63)
-where
-$t_{0}=\Bigg[\frac{\varepsilon_{3}-\mu_{1}}{2\,\varepsilon_{3}}\Bigg]a^{2}\nu^{2}_{3};\quad\quad t_{1}=\Bigg[\frac{\mu_{3}-\varepsilon_{1}}{2\,\mu_{3}}\Bigg]a^{2}\nu^{2}_{3}$ (S. 64)
-$t_{3}=\Bigg[-\frac{\varepsilon_{1}(\mu_{1}+\mu_{3})+\varepsilon_{3}(\mu_{3}-\mu_{1})}{\nu^{2}_{1}}+\frac{2\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}\Bigg]\frac{a^{2}\nu^{2}_{1}\nu^{2}_{3}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})}$
-$t_{4}=\Bigg[-\frac{\varepsilon_{3}(\mu_{1}+\mu_{3})+\varepsilon_{1}(\mu_{1}-\mu_{3})}{\nu^{2}_{1}}+\frac{2\varepsilon_{3}\mu_{3}}{\nu^{2}_{3}}\Bigg]\frac{a^{2}\nu^{2}_{1}\nu^{2}_{3}}{(\varepsilon_{1}+\varepsilon_{3})(\mu_{1}+\mu_{3})}$
-It is well-known that at large separation, the maximum contribution for the interaction energy comes from the classical zero frequency (entropic) term whose long-range interaction is
-$G^{\text{res}}_{p=0}(a,R,T)=-\frac{k_{B}T}{\pi}\int_{0}^{\infty}dk\,(t_{0}+t_{1}+t_{2}+t_{3})\,K_{0}(\nu_{3}R)$ (S. 65)
-S.4 Preliminary considerations for the impact of magnetic fluid
-In view of the complexity of the formalism, we first note some relative magnitudes of central quantities as part of the theory.
-Consider first the dispersion relation: adopting a Lorentz form for the ferrofluid magnetic susceptibility $\chi_{ms}$ with a characteristic frequency $\omega_{0m}$, we have 25
-$\mu(i\xi)=1+\frac{4\pi\chi_{ms}}{1+\xi^{2}/\omega_{0m}^{2}},$ (S. 66)
-with $\xi=-i\omega$ being the imaginary frequency.
-Here $\omega_{0m}$ is relatively low, typically $\omega_{0m}=3\times 10^{10}$ rad/s, in contrast to the characteristic dielectric $\omega_{0e}$ which is typically $\omega_{0e}=3\times 10^{16}$ rad/s.
-Hence, only zero frequency contribution is influenced strongly by the magnetic effects.
-At large separations, non-zero frequency contributions to the dispersion interaction are screened out by retardation.
-Thus, the magnetic part of the dispersion forces is an essential part of the long-range/high-temperature zero-frequency contribution.
-A characteristic property of both the non-retarded and zero-frequency entropic interaction between cylinders is its scaling $R^{-5}$ dependence32, 23, 41, 42, 31, 43, 44, 45.
-Another way of looking at the condition for nonretarded theory is to observe that the distance $R$ between the cylinders must be much smaller than the wavelength $\lambda_{0}=2\pi c/(n\omega_{0})$ corresponding to the proper frequency $\omega_{0}$ for the material ($n$ is the refractive index).
-In the electric case with $\omega_{0}=\omega_{0e}$ one gets $\lambda_{0e}\sim 50$~nm, while in the magnetic case with $\omega_{0}=\omega_{0m}$ one gets $\lambda_{0m}$ to lie in the centimeter region.
-Thus in the magnetic case the nonretarded theory is often applicable, at least when the condition $R\ll\lambda_{0}$ is satisfied.
-Also, in the absence of magnetic fields, effects due to magnetic nonlinearities are usually negligible in Casimir problems.
-However, a very recent paper discussed how the Casimir force in specific cases can be tuned by magnetic fields 37.
-The details on how we modeled the magnetic permeability of the magnetic fluid are given in the ”Materials Modeling” section.
-S.5 Material modeling
-S.5.1 Zinc oxide nanorod
-We use zinc oxide (ZnO) 34 as nanorod material as it is a wide-gap semiconductor with many advantages 46, 47, such as: ($i$) availability of large area substrates, ($ii$) high solubility of n-type dopants, (iiiiii) large exciton binding energy, ($iv$) excellent luminescent properties, ($v$) high chemical/thermal stability and radiation hardness, ($vi$) wet chemical synthesis is possible, (viivii) room-temperature ferromagnetism via Mn/Fe-doping, (viiiviii) quantum size effects on exciton and gap energies in nanocrystals, ($ix$) good UV-blocker stability stable, ($x$) low power threshold at room temperature, ($xi$) piezoelectric, (xiixii) stable polar surface, (xiiixiii) outperforming cost-effectiveness, whiteness, and UV-blocking of nanoparticles antibacterial finishing, (xivxiv) provides reinforcement in natural rubber, and ($xv$) it is bio-safe and biocompatible.
-In addition, high-quality ZnO nanorods and nanowires can be synthesized without widespread defects or dislocations in the crystal structure.
-The material properties of ZnO and magnetite (Fe3O4) were investigated within the framework of density functional theory (DFT), employing the projector augmented-wave (PAW) method as implemented in the Vienna Ab initio Simulation Package (VASP)48.
-To accurately describe the dielectric response at high energies (i.e., extending beyond 3000 Matsubara frequencies), $GW$-type core potentials were employed, an energy cutoff of 600 eV, and the initial valence configurations for the atoms are chosen to be $3s^{2}p^{6}d^{10}4s^{2}$ for Zn, $3s^{2}p^{6}d^{8}$ for Fe, and $2s^{2}p^{4}$ for O.
-The exchange-correlation interaction for the wide-gap ZnO is modeled with the hybrid functional by Heyd, Scuseria, and Ernzerhof49 with 30% Fock exchange; this yields a band gap energy of 3.34 eV.
-The hybrid functional generates proper Zn $d$-like orbitals, which is important for describing the valence $sd$-hybridization in ZnO 50.
-The irreducible Brillouin zone for the 4-atom hexagonal unit cell is sampled by a $12\times 12\times 8$ k-mesh.
-For the ferrimagnetic Fe3O4 we choose the generalized gradient approximation for solids, developed by Perdew et al.51 with an effective Hubbard correction 52 $U=5.7$ eV on the Fe $d$-like orbitals.
-A $20\times 20\times 20$ grid of k-points is used for the 14-atom face-centered cubic unit cell.
-Our calculated dielectric function of Fe3O4 agrees overall very well with the empirical spectrum by Draine and Hensley 53 in the low infrared to visible region.
-For energies below 1 eV we combine our calculated data with Draine and Hensley’s parameterised spectrum, which then accounts also for the room temperature electrical conductivity.
-From the computed electronic structures of the materials, we determine the imaginary part of the dielectric response functions in the long wavelength limit:
-$\varepsilon^{\prime\prime}(\omega)=\lim_{{\bf q}\rightarrow 0}\frac{4\pi^{2}e^{2}}{V_{\Omega}q^{2}}\sum_{v,c,{\bf k}}\delta(\epsilon_{c,{\bf k}}-\epsilon_{v,{\bf k}}-\hbar\omega)\langle u_{c,{\bf k}+{\bf e}_{\alpha}q}|u_{v,{\bf k}}\rangle\langle u_{v,{\bf k}}|u_{c,{\bf k}+{\bf e}_{\alpha}q}\rangle$ (S.
-67)
-$+\sum_{j}\frac{S_{j}\,\omega_{\rm TO,j}^{2}\Gamma_{j}\omega}{(\omega_{{\rm TO},j}^{2}-\omega^{2})^{2}+(\Gamma_{j}\omega)^{2}}+\frac{\Gamma\,\omega_{\rm pl}^{2}}{\omega(\omega^{2}+\Gamma^{2})}$
-where the first term describes the electronic response.
-Here, $V_{\Omega}$ is the unit-cell volume referenced to the three Cartesian directions ${\bf e}_{\alpha}$.
-The eigenfunctions are represented by the cell periodic part $u_{v/c}$ of the valence ($v$) or conduction ($c$) state with corresponding energy $\epsilon_{v/c,{\bf k}}$.
-The second term in Eq.
-(S.
-67) describes the ionic response, where the vibrations screen the carriers by an electric field.
-Here, $\omega_{TO}$ is the frequency of the transverse optical mode, $\Gamma_{j}$ is the damping, and $S_{j}$ is the oscillator strength of the $j$th mode in its vibration direction.
-From the density functional perturbation theory we compute the Hessian matrix of the ionic displacements to model the optical phonons.
-The third term in Eq.
-(S.
-67) describes the Drude screening with plasma frequency $\omega_{\rm pl}$ arising from nearly free carriers in the partially occupied bands.
-The dielectric functions are modeled for the imaginary frequencies and presented in Fig.
-(2) in the main text.
-S.5.2 Polystyrene nanorod and dielectric function of Toluene
-The nanoparticles we consider have large enough dimensions to approximate their effective dielectric functions with those of bulk materials.
-We use available experimental data from the literature for polystyrene (PS) nanorod33 and this reference is also used to provide the dielectric function for Toluene.
-S.6 Thermal stability
-In this section we have performed stability analysis with respect to the thermal fluctuations, following the detailed procedure outlined in Ref.21.
-Figure S.1: Colors online: Thermal stability analysis in the attractive region of total force for diameter $D=17$ nm and $\phi=5\%$.
-In this situation, the trapping width is 500 nm and force gradient in the trapping region is $7.92\times 10^{-11}\mu N$. The depth of the corresponding trapping potential is calculated and found to be
-$\Delta U\approx 4\times 10^{-23}\mathrm{J}.$
-At room temperature ($T=298.15\mathrm{K}$) the thermal energy $k_{B}T\approx 4.11\times 10^{-21}\mathrm{J}$. Comparing the two gives
-$\frac{\Delta U}{k_{B}T}\approx 0.01\Rightarrow\Delta U\ll k_{B}T.$
-As a consequence, for this specific configuration, we are in a regime where the trapping potential is much reduced compared to the thermal energy and thermal fluctuation can easily surmount the Casimir potential such that the nanoparticle remains unstable.
-This problem can be enhance using PFA theorem that is beyond our discussion.
-S.7 Numerical analysis of excited state interaction
-Safari et al.54 used perturbation theory to study dispersion interaction between a pair of chiral molecules in the presence of magnetoelectric media.
-They found strong magnetic effects.
-Many biomedical processes occur in background magnetic fluids influencing both ground and excited state interactions.
-In the long range limit the excited state dispersion interaction, as shown in Fig.
-(S.2), receives important contributions from the zero frequency term.
-Notably, excited state interactions between elongated biomolecules can be long-range and impacted by the magnetic permeability.
-Figure S.2: Colors online: Contributions to the force per unit length from an antisymmetric excited state between two cylindrical polystyrene nanoparticles with the same radius $a=20$ nm as a function of center-to-center distance $R$ when immersed in a ferrofluid with 5% concentration of magnetite nanoparticles of diameter 17 nm.
-S.8 Force calculation with gold-coated magnetite nanoparticles
-We study here the interaction forces between ZnO and PS in their ground state when suspended in a magnetic colloidal host formed by gold-coated magnetite nanoparticles dispersed in toluene.
-The objective of this study is to clarify how the compositional characteristics of magnetic nanoparticles affect the inter-nanorod potential and the resulting trapping behavior.
-Figure S.3: Colors online: The CL force per unit length between two cylindrical objects (ZnO and PS respectively) with the same radius $a=20$ nm as a function of center-to-center distance $R$ when immersed in a ferrofluid with 5% concentration of gold-coated (thickness $R_{\rm Au}=2,3,5$ nm) magnetite nanoparticles of diameter 13, 15 and 17 nm.
-We find that the trapping between ZnO and PS nanorods (for the considered cases) takes place only for magnetite nanoparticles with core size, $D=17$ nm.
-This behavior is consistent across all examined gold-layer thicknesses, namely $R_{\rm Au}=2\,\text{nm},3\,\text{nm}$, and 5 nm in Fig.
-(S.3).
-We find this trapping to be thermally unstable which we discuss in Sec.
-(S.6) and can be overcome via PFA theorem.
-Data availability: The code and data that support the findings of this study are available in Github.